### v1

# Large-scale, Independent and Comprehensive study of the power of LLMs for test case generation

Abstract.
Unit testing, crucial for identifying bugs in code modules like classes and methods, is often neglected by developers due to time constraints. Automated test generation techniques have emerged to address this, but often lack readability and require developer intervention. Large Language Models (LLMs), like GPT and Mistral, show promise in software engineering, including in test generation. However, their effectiveness remains unclear.
This study conducts the first comprehensive investigation of LLMs, evaluating the effectiveness of four LLMs and five prompt engineering techniques, for unit test generation. We analyze 216 300 tests generated by the selected advanced instruct-tuned LLMs for 690 Java classes collected from diverse datasets. We assess correctness, understandability, coverage, and bug detection capabilities of LLM-generated tests, comparing them to EvoSuite, a popular automated testing tool. While LLMs show potential, improvements in test correctness are necessary. This study reveals the strengths and limitations of LLMs compared to traditional methods, paving the way for further research on LLMs in software engineering.
1. Introduction
Unit tests plays a crucial role in software development, as they validate individual functions or code units, ensuring their correctness and reliability (Beck, 2000; Siddiqui, 2021; Shore and Warden, 2021). In addition to verification, unit tests help in early bug detection, enhance code quality, provide documentation, support debugging, and reduce the risk of costly errors in production. However, manual unit test creation can be arduous and discouraging, leading some developers to skip testing entirely (Daka and Fraser, 2014; Kochhar et al., 2013b, a).
To cope with the challenges of test creation by developers, automated test generation techniques have gained traction, including methods like fuzzing (Miller et al., 1990; Zeller et al., 2019), feedback-directed random test generation (Toffola et al., 2017; Csallner and Smaragdakis, 2004; Pacheco and Ernst, 2007; Pacheco et al., 2008; Selakovic et al., 2018; Arteca et al., 2022), dynamic symbolic execution (Godefroid et al., 2005; Sen et al., 2005; Cadar et al., 2008; Tillmann et al., 2014), and search-based and evolutionary approaches (Fraser and Arcuri, 2011a, b). These techniques aim to create tests that maximize coverage through static or dynamic analyses. EvoSuite (Fraser and Arcuri, 2011b), a prominent search-based method, uses evolutionary algorithms to generate test suites for Java classes, aiming to maximize coverage. However, despite its effectiveness, the generated tests often lack readability (Almasi et al., 2017; Grano et al., 2018; Roy et al., 2020), contain obscure variable names (Daka et al., 2015), lack assertions (Panichella et al., 2020) or include overly general or excessive assertions (Panichella et al., 2020), and have shortcomings in detecting real bugs (Bacchelli et al., 2008; Pinto and Vergilio, 2010; Shamshiri et al., 2015; Almasi et al., 2017). Although these tests can inspire manual development of high-coverage test suites, they often lack naturalness and practical usability (Yuan et al., 2023; Shin et al., 2023; Almasi et al., 2017; Panichella et al., 2020; Alagarsamy et al., 2023).
Given the limitations of traditional analysis-based test generation methods, recent research has focused on using advanced deep learning techniques to improve unit test generation. Some techniques treat test generation as a neural machine translation problem, fine-tuning models to translate classes or methods into corresponding test prefixes or assertions (Tufano et al., 2020; Watson et al., 2020; Dinella et al., 2022; Nie et al., 2023; Shin et al., 2023; Alagarsamy et al., 2023). Recently, Large Language Models (LLMs), such as GPT-3 (Brown et al., 2020), have been leveraged for this purpose due to their extensive parameter sets and versatility (Fan et al., 2023). Studies (Wang et al., 2024; Fan et al., 2023; Kang et al., 2023; Lemieux et al., 2023; Rao et al., 2023; Chen et al., 2023; Siddiq et al., 2023) have then shown that LLMs can significantly enhance the quality of automated test case generation, streamlining the process for developers. These LLM-based test generation approaches are claimed to address issues of readability, understandability, and meaningfulness (Siddiq et al., 2024b; Yuan et al., 2023; Tang et al., 2024). The emergence of LLMs therefore may mark a significant evolution in software development, enabling more automated, intelligent, and efficient testing processes.
Despite the potential of advanced LLMs in software engineering tasks, it is crucial to critically evaluate their actual effectiveness in generating unit test cases. We are concerned in several main dimensions:
Data leakage: Sallou et al. (Sallou et al., 2023) caution that current methods might miss critical information and use potentially irrelevant metrics, with data leakage risks leading to overly optimistic results. There is thus a need to assess the capability of LLMs with ”unseen” code for test generation.
LLM diversity: The promise of LLMs for test generation has been mostly studied based on the capabilities of GPT-3, which is a closed model. Other models, such as Mistral 7B (Jiang et al., 2023), may be relevant to assess for this daily task in software engineering.
Prompting strategy: A recent survey by Wang et al. (Wang et al., 2024) has revealed that Zero-shot and Few-shot learning are the two most applied prompting techniques for LLM-based test case generation. Yet, other prompting strategies (Amatriain, 2024), such as Chain-of-Thought (Wei et al., 2022) and Tree-of-Thoughts (Yao et al., 2024), have already shown to substantially improve results in various software engineering tasks.
Evaluation methodology: Beyond classical metrics such as code coverage, few works actually attempt to assess whether the LLM-generated test cases are readable or actually contribute to bug detection.
This paper: Our work investigates the effectiveness of LLMs in unit test generation and how prompt engineering techniques might affect their performance. Drawing inspiration from (1) works that compare Search-Based Software Testing (SBST) against ChatGPT (Tang et al., 2024), (2) comprehensive surveys in the field (Wang et al., 2024), and (3) insightful early experiments on the threats of using LLMs in software engineering (Sallou et al., 2023), we undertake a large-scale study to evaluate the potential of LLMs and prompt engineering to improve unit test case generation.
Unlike much of the prior research, which focus mainly on criteria related to syntax accuracy, compilability, and at best, code coverage, our study further evaluates these criteria along with the capability of LLMs to generate functional test suites that can identify real-world software bugs.
We assess four popular LLMs, GPT (3.5 and 4) from OpenAI and Mistral (7b and 8X7B) from MistralAI, and employ five prompt techniques implemented in a way that is suitable for test generation. Our comparative analysis examines the unit test suites generated by these LLMs against those from EvoSuite, a widely-known SBST tool. We rely for our experiments on two well-known datasets, SF110 (Fraser and Arcuri, 2014) and Defects4J (Just et al., 2014), as well as a new dataset, collected based on the criterion111This Custom dataset includes projects released on GitHub after May 2023, to ensure the study’s findings are pertinent and free from biases associated with pre-existing LLM training data. that the test cases must have been “unseen” by the LLMs during their training.
Our main contributions are as follows:
We provide insights on the capabilities of LLMs for unit test generation based on a comprehensive analysis of the impact of prompt engineering. To the best of our knowledge, this constitutes the first work in the literature that addresses this question.
We describe our findings based on a comprehensive array of evaluation metrics, including correctness, readability, code coverage, and the capacity for bug detection. By leveraging various datasets and models, we provide a robust basis for comparisons and future research. In particular, our assessment on an “unseen” dataset strengthens our study’s relevance and mitigates biases related to data leakage in LLMs.
We provide a comparative evaluation of LLMs against a traditional SBST method.
To aid ongoing and future research in this field, we release a comprehensive benchmark including (1) a labeled dataset of 216,300 tests generated by four instruction-tuned LLMs across 690 Java classes, (2) and performance evaluation scripts. These artifacts are available at:
The remainder of this paper is structured as follows: Section 2 details the background of LLMs for unit test generation, and prompt engineering. Section 3 outlines our experimental setup, followed by Section 4, which discusses the evaluation results. While Section 5 addresses potential study threats and limitations, Section 6 reviews related work, and Section 7 offers concluding remarks.
2. Background
In this section, we introduce the fundamental concepts essential for understanding our paper. We succinctly cover LLMs for unit test generation along with prompt engineering, which constitutes the central ideas of this study.
2.1. LLM-based Unit Test Generation
Large language models (LLMs) are extensive neural networks built primarily on the Transformer architecture (Vaswani et al., 2017) and trained on vast text corpora and code using self-supervised techniques like Masked Language Modeling and Causal Language Modeling (Shanahan, 2024; Naveed et al., 2023; Hadi et al., 2023; Shanahan, 2024). These models, including encoder-only like BERT (Devlin et al., 2018), decoder-only like GPT (Brown et al., 2020; Yenduri et al., 2024), and encoder-decoder models like T5 (Raffel et al., 2020), are often enhanced with methods like Reinforcement Learning from Human Feedback (RLHF) (Ziegler et al., 2019; Bai et al., 2022) to better align with human intentions. When interacting with these models, the instructions provided alongside any input data form what is referred to as a prompt (Siddiq et al., 2024b; Amatriain, 2024; Vogelsang and Fischbach, 2024; Siddiq et al., 2024a). Indeed, a prompt is a text description in natural language that specifies the task to be performed by an LLM. As shown in Figure 1, an LLM can be prompted 222https://chatgpt.com/share/99d6468e-98ed-40b1-9b14-17bb9c239f9c to generate a test suite for a Divider Class, and it will respond by producing the test suite, illustrating its ability to understand and execute the task specified in the prompt.
LLMs have been notably applied in various fields, with several specialized models tailored for software engineering tasks like code synthesis and test generation (Fan et al., 2023; Wang et al., 2024). Prominent examples, such as Codex (Chen et al., 2021) and StarCoder (Li et al., 2023), can perform test generation. Despite the widespread usage of LLMs in code and test generation tasks (Wang et al., 2024; Fan et al., 2023; Yuan et al., 2023; Siddiq et al., 2024b; Schäfer et al., 2023; Chen et al., 2023; Tang et al., 2024), their effectiveness in unit test generation through prompt engineering remains under-explored. This study addresses this gap by evaluating the capabilities of LLMs in generating unit tests, marking a significant exploration into their application in software engineering.
2.2. Prompt Engineering
Prompt engineering involves creating prompts which enable LLMs to effectively address specific tasks (Amatriain, 2024; Siddiq et al., 2024a; Vogelsang and Fischbach, 2024; Naveed et al., 2023; Sahoo et al., 2024). This technique is particularly useful in complex problem-solving scenarios, where a single line of reasoning might not be enough (Amatriain, 2024; Kojima et al., 2022; Wei et al., 2022; Reynolds and McDonell, 2021). It allows LLMs to adopt a more human-like approach by considering multiple possibilities before reaching a conclusion. This emerging field involves crafting prompts that guide LLMs to generate desired outputs, ranging from broad descriptions to detailed specifications. The importance of prompt engineering is twofold: Firstly, highly specific prompts yield precise responses, while more general prompts may produce less relevant outputs (Si et al., 2022). Secondly, adjusting prompt instructions can help control the output format of LLMs, minimizing the need for post-processing and improving performance evaluation (Macedo et al., 2024).
Various studies (Fan et al., 2023; Wang et al., 2024) have leveraged LLMs and prompt engineering for automated software engineering tasks, commonly using basic approaches such as Zero-shot (Kojima et al., 2022) and Few-shot learning (Brown et al., 2020). Recently, research has broadened prompt engineering to include techniques such as Chain-of-Thought (Wei et al., 2022), self-consistency (Wang et al., 2022), and Tree-of-Thoughts (Wang et al., 2022). Figure 2 illustrates various examples of prompts (additional examples of prompt engineering techniques are available in the replication package).
Zero-shot learning enables models to generalize to new tasks without specific examples but may falter with complex tasks (Sahoo et al., 2024; Kojima et al., 2022). Few-shot learning, or in-context learning, uses a few examples to improve model performance on novel tasks (Brown et al., 2020; Reynolds and McDonell, 2021). Chain-of-Thought (CoT), introduced by Wei et al. (Wei et al., 2022), outlines intermediate reasoning steps and can be enhanced by combining it with few-shot prompts for complex reasoning. Self-consistency, proposed by Wang et al. (Wang et al., 2022), uses diverse reasoning paths to select the most consistent answers, improving deep reasoning tasks. Tree-of-Thoughts (ToT), a newer technique proposed by Yao et al. (Wang et al., 2022) and (Long, 2023), builds on Chain-of-Thought by employing a systematic exploration through a ”tree” of reasoning steps, integrating search algorithms for more strategic problem-solving. This study explores various prompt engineering methods to adapt pre-trained LLMs for unit test generation, emphasizing their application in guiding LLM behavior in test generation tasks.
3. Experimental Setup
3.1. Overview
Figure 3 outlines our methodology, detailing the steps from prompting LLMs using various datasets to generate unit test cases, through to evaluating these tests. The process involves generating test suites for each class, repeated thirty times. Syntactically correct test suites are isolated, compiled, and validated, discarding any incorrect or textual outputs. The first two research questions focus on analyzing LLM-generated tests in terms of output format, syntax correction, compilability (RQ1), and comprehensibility (RQ2). The last two questions compare the performance of LLM-generated test suites with those created by EvoSuite in terms of code coverage (RQ3) and bug detection (RQ4).
3.2. Research Questions
This study evaluates the effectiveness of LLMs in test case generation and examines the impact of prompt engineering on their performance. It includes an investigation structured around four specific Research Questions (RQs).
RQ1: How does prompt engineering influence the generation of syntactically correct and compilable test suites by LLMs? We investigate the potential for suitable prompt engineering to guide LLMs towards producing test suites that: ❶ are syntactically correct: this ensures the code adheres to the programming language’s rules, increasing the likelihood of successful compilation; and ❷ compile without errors: compiled code enables execution but doesn’t guarantee functionality. Findings in the study of this research will contribute to improving the ability of LLMs to generate well-formed test suites that can be automatically extracted for code analysis, ultimately promoting development efficiency.
RQ2: Beyond syntactic correctness, how well do LLMs generate human-readable and maintainable test suites? This research question examines the readability and maintainability of test suites produced by LLMs. Well-written test suites are essential for developers to: ❶ Understand the test logic: Clear and concise code makes it easier to grasp the purpose and functionality of the tests; ❷ Maintain the tests over time: As code evolves, readable tests require less effort to update and keep current; and ❸ Integrate tests into existing codebases: Easy-to-understand tests facilitate smooth merging with existing code. By evaluating these aspects in this research question, we aim to determine whether LLM-generated test suites are practical for real-world development workflows.
RQ3: How does the code coverage achieved by LLM-generated test suites compare to those generated by a Search-Based Software Testing technique? This research question investigates the effectiveness of LLMs in generating test suites that thoroughly cover the code under test. Code coverage is a key metric indicating the percentage of code executed by the test suite. Higher coverage suggests a more comprehensive evaluation, potentially leading to better bug detection. EvoSuite, a prominent SBST tool, serves as a benchmark to assess the capabilities of LLMs in this domain.
RQ4: How effective are LLM-generated test suites in detecting bugs compared to EvoSuite-generated test suites? This research question examines the practical value of LLM-generated test suites. We compare their ability to uncover software bugs against the ability of EvoSuite, a well-established SBST technique. High code coverage, which is a known strength of SBST techniques, is crucial, but it doesn’t guarantee efficient bug detection. Therefore, with this research, we aim to identify if LLMs offer any strengths in finding software bugs beyond the capabilities of traditional SBST techniques.
3.3. Large Language Models
Our study focuses on four instruction-tuned decoder-only models known for their enhanced ability to follow instructions compared to base models. These models belong to the OpenAI and MistralAI families and are detailed in Table 1.
One such model, GPT-3.5-turbo, has a token limit of 4,096. Research suggests that exceeding this limit when splitting a class into test case generation methods can hinder the model’s understanding of the class context, potentially reducing test quality (Tang et al., 2024). To ensure a fair comparison, all models in our study are restricted to this token limit, and any inputs exceeding it are excluded. Token counts are determined using OpenAI’s tiktoken333https://github.com/openai/tiktoken library.
Furthermore, all models are set to a standard temperature of 0.7. Our selection criteria prioritized models that excel in code generation and possess accessible APIs, making them suitable for both academic research and industrial applications (Wang et al., 2024; Fan et al., 2023).
| Dataset | Description | Number of Projects | Number of Classes | Number of Bugs | Complexity Overview | ||
| Low | Moderate | High | |||||
| SF110 | Derived from DynaMOSA benchmark (Panichella et al., 2017), includes 346 Java classes from 117 projects. | 74 | 182 | N/A | 170 | 138 | 120 |
| Defects4J | Contains 835 reproducible bugs across 17 open-source projects, focuses on detailed test suites to expose bugs. | 17 | 477 | 403 | 254 | 177 | 130 |
| Custom mini Dataset (CMD) | Includes ODC a collaborative project to enhance database functionalities and Conductor OSS an open-source project management tool aimed at efficiency and scalability. Combines bugged and fixed classes for bug detection analysis. | 2 | 31 | 20 | 16 | 9 | 5 |
3.4. Prompt Design
Unit test generation research highlights the importance of practical and focused methods that avoid overly intricate solutions (e.g., (Siddiq et al., 2024b; Yuan et al., 2023; Chen et al., 2023; Tang et al., 2024)). To achieve this, we designed prompts with varying complexity levels based on a clear and well-structured template. Each prompt consists of two parts: a Natural Language Description (NLD) outlining the task and a Source Code placeholder to insert the Class Under Test (CUT). We will now explore the details of each part, explaining how they align with established unit testing practices.
Natural Language Description: Drawing on established best practices for prompt effectiveness, we designed our prompts as Natural Language Descriptions (NLDs) that outline the task for LLMs. These NLDs incorporate three key elements: (i) a role-playing directive to mimic a professional software tester (or several testers), which is known to improve test generation (Tang et al., 2024; Chen et al., 2023); (ii) a task-specific directive to generate a complete Junit 4 test file for a specified class, and (iii) a control statement requiring the output to be enclosed between specific start and end markers, ensuring structured output.
Source Code: Consistent with existing studies on learning-based unit test generation (Tufano et al., 2020), our prompts vary based on the engineering technique, either including (i) just the class name and its source code (Figure 2(a)), or (ii) the class name, its source code, plus an example of another class’s source code and its test suite, to provide models with context or comparative examples (Figure 2(b)).
This study investigates the impact of varying prompt complexity on the quality of generated test cases. We employ five distinct prompt engineering techniques: Zero-shot Learning (ZSL), Few-shot learning (FSL), Chain-of-Thought (CoT), Tree-of-thoughts (ToT), and a novel technique called Guided Tree-of-thoughts (GToT). ZSL provides prompts without specific examples, forcing models to generalize to new tasks. FSL offers a few examples to guide the model towards better performance. CoT structures prompts to encourage logical reasoning, while ToT promotes exploration of multiple reasoning paths for more informed decisions. GToT attempts to exploit advantages of both CoT and ToT by guiding the reasoning of each solicited expert within the prompt itself. This research focuses on these techniques as test case generation demands a deep understanding of the code and the ability to reason through complex scenarios. Further details and comparisons of these prompt templates are available in the replication package.
3.5. The EvoSuite SBST tool
EvoSuite Configuration
This section details the configuration used for EvoSuite, a well-established Search-Based Software Testing (SBST) tool, in our study. We employed EvoSuite with its default settings, including the DynaMOSA algorithm (Panichella et al., 2017). This algorithm focuses on achieving high coverage for branches, methods, and lines of code within the test suite.
To account for the inherent randomness of genetic algorithms used by EvoSuite, we executed the tool 30 times for each class under test. This aligns with previous research suggesting a three-minute time budget per test case generation run (Fraser et al., 2015; Arcuri and Fraser, 2013; Shamshiri et al., 2015). This time limit reflects practical considerations faced by developers during the testing process.
Hardware
For reference, EvoSuite experiments were conducted on an AMD EPYC 7552 48-Core Processor running at 3.2 GHz with 640 GB of RAM. However, it’s important to note that not all available resources were necessarily utilized during the tests.
3.6. Datasets
To minimize selection bias and ensure generalizability, we employed three diverse datasets: SF110, Defects4J, and a Custom Mini Dataset (CMD). Both SF110 and Defects4J were used to evaluate both EvoSuite and LLM performance. The CMD, containing GitHub projects created after May 2023, was exclusively tested with LLMs due to compatibility issues with EvoSuite’s Java version requirements (see Issue #433 on the EvoSuite GitHub repository(evo, [n. d.])). This dataset focused on well-maintained projects from the OceanBase Developer Center (ODC) (OceanBase, 2023) and Conductor OSS (conductor oss, 2023).
The datasets encompass projects with code of varying levels of cyclomatic complexity (detailed in Table 2), ranging from low to high. This diversity helps mitigate potential overfitting on specific complexity profiles and provides valuable insights into the potential of LLMs for real-world automated test case generation.
Notice. Due to budgetary reasons, we only run GPT-4 and Mixtral-8x7B on the smallest dataset (CMD). Other models (GPT-3.5-Turbo and Mistral-7B) are run on all datasets.
4. Experimental Results
| Model Name | # of Generated Files | # of Files with code | MSR | CSR | ||||||||
| SF110 | Defects4J | CMD | SF110 | Defects4J | CMD | SF110 | Defects4J | CMD | SF110 | Defects4J | CMD | |
| GPT 3.5-turbo | 27 300 | 71 550 | 4 650 | 23 048 | 60 565 | 4 633 | 84.42% | 84.65% | 99.64% | 78.10% | 78.34% | 90.79% |
| GPT 4∗1 | – | – | 4 650 | – | – | 4 356 | – | – | 93.68% | – | – | 83.59% |
| Mistral 7B | 27 300 | 71 550 | 4 650 | 25 664 | 65 180 | 4 226 | 94.00% | 91.10% | 90.88% | 79.97% | 80.35% | 83.61% |
| Mixtral 8x7B∗1 | – | – | 4 650 | – | – | 2 947 | – | – | 63.38% | – | – | 53.57% |
| Total/Average (%) | 54 600 | 143 100 | 18 600 | 48 712 | 125 745 | 16 162 | 89.21 | 87.88 | 86.90 | 79.04 | 79.35 | 77.89 |
1
∗Note that due to budget constraints, we only paid to query GPT-4 and Mixtral 8x7B for the small dataset (CMD).
| Model Name | Syntactically Correct/Compilable Code SF110 (%) | Syntactically Correct/Compilable Code Defects4J (%) | Syntactically Correct/Compilable Code CMD (%) | ||||||||||||
| ZSL | FSL | CoT | ToT | GToT | ZSL | FSL | CoT | ToT | GToT | ZSL | FSL | CoT | ToT | GToT | |
| GPT 3.5-turbo | 21.30/5.94 | 17.30/4.21 | 18.05/5.46 | 19.70/4.08 | 23.65/9.67 | 21.51/2.95 | 16.43/2.22 | 17.51/3.23 | 20.02/2.26 | 24.54/6.76 | 21.70/2.70 | 20.77/2.44 | 20.09/5.92 | 16.39/2.68 | 21.06/3.46 |
| GPT 4 | –/– | –/– | –/– | –/– | –/– | –/– | –/– | –/– | –/– | –/– | 23.00/1.78 | 22.90/2.21 | 20.68/2.65 | 17.60/1.90 | 15.82/2.39 |
| Mistral 7B | 21.25/1.09 | 21.15/1.04 | 20.95/0.91 | 20.43/1.74 | 16.23/3.37 | 20.26/0.45 | 20.23/0.43 | 20.65/0.27 | 20.13/0.44 | 18.73/1.09 | 20.37/0.00 | 20.29/0.03 | 20.60/0.00 | 20.11/0.03 | 18.62/0.31 |
| Mixtral 8x7B | –/– | –/– | –/– | –/– | –/– | –/– | –/– | –/– | –/– | –/– | 18.85/0.16 | 23.36/0.12 | 18.51/0.08 | 15.58/0.44 | 23.60/0.16 |
| Total/Average (%) | 21.27/3.52 | 19.22/2.62 | 19.50/3.19 | 20.06/2.91 | 19.94/6.52 | 20.89/1.70 | 18.33/1.32 | 19.08/1.75 | 20.08/1.35 | 21.63/3.93 | 21/1.16 | 21.83/1.20 | 19.97/2.16 | 17.42/1.26 | 19.78/1.58 |
1
∗ZSL: Zero-shot learning, FSL: Few-shot learning, CoT: Chain-of-Thought, ToT: Tree-of-Thoughts, GToT: Guided Tree-of-thoughts.
4.1. [RQ1]: Output Format, Syntactical Correctness, and Compilability.
[Experiment Goal]: We aim to explore how different prompts may lead the LLMs to exhibit difference effectiveness in producing syntactically correct and compilable test suites in the required output format.
[Experiment Design]: We consider five prompt designs to instruct the LLMs to generate test cases and use regular expressions to extract the source code of test suites from the LLM outputs. To evaluate the effectiveness of the techniques, we adopted the Match Success Rate (MSR) and Code Extraction Success Rate (CSR) metrics introduced by Macedo et al. (Macedo et al., 2024). MSR measures the proportion of outputs where the output could be matched by a pre-defined regexp pattern. It is thus a proxy to measure to what extent the LLM adheres to the format specified in the prompt. CSR computes the proportion of cases where a source code can be extracted within the matched output.
We analyze the extracted code using Javalang 0.13.0 (c2nes, 2012) to verify its syntactical correctness. If syntactically correct, the test suite then undergoes testing on the Java Virtual Machine (JVM) to check for compilability (for this RQ) as well as executability, and any runtime errors (in subsequent RQs). We also assess the quality of the generated test cases using SpotBugs 4.8.3 (spotbugs, 2016), an open-source static analyzer. SpotBugs checks Java programs for over 400 types of bugs and suboptimal coding practices. This analysis identifies how well the test cases detect common programming mistakes and ensure adherence to coding best practices.
[Experiment Results]: Our study generated 216 300 files, with 190 619 containing source code and 171 187 cases where syntactically correct test suites are provided. Table 3 provides performance measurements of LLMs in terms of MSR and CSR.
GPT 3.5-turbo achieves the highest MSR (99.64%) and CSR (90.79%) on the Custom dataset. Mistral 7B also achieve high performance on the SF110 and Defects4J datasets. In contrast, Mixtral 8x7B struggles on the Custom dataset, in comparison to other models. When using GPT-4, we record a high CSR score at 83.59%.
Table 4 summarizes the performance of LLMs in generating syntactically correct (i.e., parsable) and compilable test suites. Overall, the metrics clearly show that LLMs underperform in generating syntactically correct test cases and are further struggling to generate compilable ones. In the best case scenario, GPT-3.5-Turbo with GToT prompting offers the highest effectiveness score on the Defects4J dataset with only 24.54% of generated test cases being syntactically correct, and 6.76% of these being also compilable. Overall, the results show that, on average, only 12% of the syntactically correct test suites compile successfully.
Following compilation, we perform static analysis using SpotBugs, which identifies 8 461 potential issues across 20 530 test cases. We classify these issues (aka bugs) into four severity levels: Scariest (most critical, affecting functionality and security), Scary (significant issues if unaddressed), Troubling (minor issues), and Of Concern (indicating lower severity). The majority (55.58%) of the issues are labeled as only Of Concern, i.e., they are informational and of minimal risk. We find that Few-shot prompts lead the LLMs to generate most (61.72%) of the ’Scariest’ bugs. We nevertheless note that GToT is associated with most of the ’Scary’ (63.92%), ’Troubling’ (69.10%), and ’Of Concern’ (33.77%) bugs. Table 5 details the statistics of issues found by SpotBugs in the test suites generated by GPT-3.5-turbo using varying prompts. ToT generates the least issues while CoT generates only ’of concern’ issues. In contrast, FSL leads to a large number of test cases with ’Scariest’ issues.
| Rank-Level | ZSL | FSL | CoT | ToT | GToT | Total |
| Of Concern | 844 | 619 | 794 | 51 | 1106 | 3414 |
| Scariest | 0 | 587 | 0 | 1 | 0 | 588 |
| Scary | 0 | 0 | 0 | 2 | 825 | 827 |
| Troubling | 227 | 14 | 0 | 3 | 825 | 1069 |
| Model | Dataset | Prompt | Classes Affected | |||
| Of Concern | Scariest | Scary | Troubling | |||
| GPT 3 | SF110 | ZSL | 600 | 0 | 0 | 0 |
| FSL | 410 | 410 | 0 | 0 | ||
| CoT | 553 | 0 | 0 | 0 | ||
| ToT | 20 | 0 | 1 | 0 | ||
| GToT | 825 | 0 | 825 | 825 | ||
| Defects4J | ZSL | 212 | 0 | 0 | 212 | |
| FSL | 177 | 177 | 0 | 0 | ||
| CoT | 219 | 0 | 0 | 0 | ||
| ToT | 21 | 1 | 1 | 2 | ||
| GToT | 242 | 0 | 0 | 0 | ||
| CMD | ZSL | 17 | 0 | 0 | 0 | |
| FSL | 18 | 0 | 0 | 0 | ||
| CoT | 22 | 0 | 0 | 0 | ||
| ToT | 3 | 0 | 0 | 0 | ||
| GToT | 20 | 0 | 0 | 0 | ||
| GPT 4 | CMD | ZSL | 15 | 0 | 0 | 15 |
| FSL | 14 | 0 | 0 | 14 | ||
| CoT | 0 | 0 | 0 | 0 | ||
| ToT | 7 | 0 | 0 | 1 | ||
| GToT | 19 | 0 | 0 | 0 | ||
| Mistral7B | SF110 | ZSL | 129 | 129 | 129 | 0 |
| FSL | 135 | 135 | 0 | 0 | ||
| CoT | 125 | 0 | 125 | 125 | ||
| ToT | 212 | 212 | 212 | 0 | ||
| GToT | 347 | 0 | 0 | 0 | ||
| Defects4J | ZSL | 48 | 48 | 0 | 0 | |
| FSL | 60 | 60 | 0 | 0 | ||
| CoT | 32 | 32 | 0 | 0 | ||
| ToT | 63 | 63 | 0 | 0 | ||
| GToT | 131 | 0 | 0 | 0 | ||
| CMD | ZSL | 0 | 0 | 0 | 0 | |
| FSL | 0 | 0 | 0 | 0 | ||
| CoT | 0 | 0 | 0 | 0 | ||
| ToT | 1 | 0 | 0 | 0 | ||
| GToT | 0 | 0 | 0 | 0 | ||
| Mixtral8x7B | CMD | ZSL | 2 | 0 | 0 | 0 |
| FSL | 0 | 0 | 0 | 0 | ||
| CoT | 0 | 0 | 0 | 0 | ||
| ToT | 0 | 0 | 0 | 0 | ||
| GToT | 4 | 0 | 0 | 0 | ||
1
∗ ZSL: Zero-shot Learning, FSL: Few-shot Learning, CoT: Chain-of-Thought, ToT: Tree-of-Thoughts, GToT: Guided Tree-of-Thoughts.
4.2. [RQ2]: Understandability.
[Experiment Goal]: We aim to assess the readability of code generated by LLMs through code style checking and understandability measures.
[Experiment Design]: We use Checkstyle 10.14.2 (checkstyle, 2001) to ensure compliance with Sun Code Conventions (Oracle, 1999) and Google Java Style (Google, 2015), verifying that LLM-generated test suites adhere to these standards. Additionally, we use metrics from Dantas et al. (Dantas and Maia, 2021) for evaluating cognitive and cyclomatic complexities with the PMD 7.0.0-rca tool (PMD, 2012). Cyclomatic complexity quantifies independent paths in the code, suggesting its complexity and error potential, while cognitive complexity focuses on the code’s ease of understanding, considering factors like structure and naming. These measures are essential for evaluating code maintainability and pinpointing potentially confusing or complex parts. The default thresholds in PMD for cyclomatic and cognitive complexity are 10 and 15, respectively.
Code Style Checking Results. Figure 4 visualizes the most common violations for each prompting technique according to both Sun and Google standards.
– Google Java Style: For Google coding standards, FSL achieved the best adherence with the fewest violations, followed by CoT. Other techniques showed lower adherence, reducing code readability. Common issues (Figure 4(a)) included incorrect indentation, improper import order, missing Javadoc documentation, wildcard imports, and long lines.
– Sun Code Conventions: For Sun coding standards, Few-shot learning was again the most compliant, followed by CoT and Zero-shot learning. ToT and GToT had higher violations, indicating weaker adherence and readability. Common issues (Figure 4(b)) included inadequate method documentation, use of ”magic” numbers, regex problems, excessive line length, and insufficient package documentation. These findings highlight the impact on code maintainability and clarity.
These evaluations help gauge the frequency and severity of style violations, with fewer average violations indicating better compliance with coding standards and improved readability. Detailed results for Checkstyle violations against Google and Sun Code Styles are documented in the replication package, along with distribution plots and statistical measures used to compare prompt strategies against coding standards.
Code Understanding Results. Table 7 shows the complexity Results by Prompt Engineering Techniques. Cyclomatic complexity helps assess the maintainability of the codebase, with higher values indicating lower maintainability, increased risk of errors, and greater difficulty in understanding and modifying the code. Lower cognitive complexity indicates more readable and understandable code. Note that PMD and Sonar, each with different levels of cyclomatic complexity. To facilitate comparison, we selected PMD levels but adjusted them according to Sonar’s standards (Campbell, 2023).
| Cyclomatic Complexity Level | ZSL | FSL | CoT | ToT | GToT |
| Low complexity (1-4) | 3 601 | 2 794 | 3 669 | 2 960 | 7 468 |
| Moderate complexity (5-7) | 4 | 11 | 6 | 6 | 1 |
| High complexity (8-10) | 1 | 3 | 4 | 2 | 0 |
| Very High complexity (11+) | 0 | 0 | 0 | 0 | 0 |
| Cognitive Complexity Level | ZSL | FSL | CoT | ToT | GToT |
| Low complexity (¡5) | 3 600 | 2 804 | 3 674 | 2 967 | 7 469 |
| Moderate complexity (5-10) | 4 | 4 | 5 | 1 | 0 |
| High complexity (11-20) | 1 | 0 | 0 | 0 | 0 |
| Very High complexity (21+) | 1 | 0 | 0 | 0 | 0 |
1
ZSL: Zero-shot learning, FSL: Few-shot learning, CoT: Chain-of-Thought, ToT: Tree-of-Thoughts, GToT: Guided Tree-of-Thoughts.
– Cognitive Complexity: Most test suites generated exhibit low complexity (Table 7(b)), particularly those produced by the Guided Tree-of-thoughts (GToT) technique, which produced the most low complexity cases. This demonstrates GToT’s ability to generate simple and maintainable code. Instances of moderate to very high complexity are rare, indicating the overall effectiveness of these techniques in producing straightforward and understandable code.
– Cyclomatic Complexity: Most generated test suites, especially with the GToT technique, exhibit low complexity (Table 7(a)), indicating effective, maintainable code creation. Moderate and high complexity instances are present but don’t reach very high levels, ensuring good maintainability. Other techniques also produce manageable code but may occasionally result in higher complexity, impacting maintainability and comprehension.
4.3. [RQ3]: Code Coverage.
[Experiment Goal]: We aim to evaluate how LLM-generated test suites fair against SBST-generated test suites in terms of code coverage.
[Experiment Design]: We consider EvoSuite as the baseline SBST tool for comparison. We measure line, instruction, and method coverage using the JaCoCo 0.8.8 tool (JaCoCo, 2022). For each class, we generate the test suites by each different LLM/prompt combination and Evosuite.
[Experiment Results]: We assess the effectiveness of each model/prompt combination or technique by employing statistical measures, such as sample size, minimum, maximum, quartiles, median, and mean, to compare code coverage metrics. Due to space constraints, all detailed results are provided in the replication package. Figure 5 illustrates example comparison results for Line coverage for Mistral-7B, GPT-3.5-Turbo and EvoSuite. We recall that since EvoSuite is incompatible with the Custom dataset (cf. Section 3.6), it was not applied for this dataset.
LLMs exhibit variable performance across different metrics, particularly GPT-3.5 Turbo, which is heavily influenced by advanced prompting techniques like GToT and CoT (Figure 5). In the Custom dataset, GPT-3.5 Turbo achieves near-perfect median coverage (100%), suggesting effective adaptation to novel software contexts. Interestingly, it outperforms GPT-4 in this dataset, potentially due to the smaller size or limitations of GPT-4 for this specific task. However, GPT-3.5 Turbo’s performance varies across other datasets. It demonstrates a low median line coverage of 3.57% in SF110 but achieves significantly higher coverage in Defects4J (82.74% lines, 85.0% instructions, 100% methods) compared to SF110 (2.0% lines, 16.67% instructions). Mistral 7B also shows inconsistent performance, particularly struggling in SF110 with low median coverages (3.03% lines, 2.0% instructions, 20.0% methods). However, it achieves improved results in Defects4J, with higher median coverages (63.4% lines, 66.0% instructions, 100% methods).
Performance discrepancies between SF110 and Defects4J suggest the influence of dataset complexity or potential data leakage (discussed in Sallou et al. (Sallou et al., 2023)). Table 2 supports this notion, as SF110 exhibits the highest proportion of classes in moderate and high complexity categories despite its smaller sample size. Limited accessibility of SF110, unlike Defects4J’s broad availability (potentially inflating metrics as discussed in (Sallou et al., 2023)), might mitigate leakage risks. Favorable results in the new, smaller dataset could be skewed by its lower variability and simplicity. The dominance of low complexity classes in this dataset (Table 2) suggests potentially simpler and more modularized code, which may be easier to understand and less challenging for the models. EvoSuite consistently achieves near-perfect coverage across all metrics, demonstrating its effectiveness as a search-based software testing tool. In contrast, LLMs (e.g., GPT-3.5 Turbo and Mistral 7B) cover significant code portions but fall short of EvoSuite’s comprehensive levels. This stems from their differing test generation approaches: EvoSuite prioritizes maximizing coverage, while LLMs rely on learned patterns and instructions, which may not always guarantee thorough coverage.
Our findings highlight the importance of refining prompt engineering to fully leverage LLM capabilities across diverse testing scenarios. Improved prompting, combined with exploring SBST methodologies, shows potential for enhancing test suite quality and effectiveness. Additionally, the superior performance of GPT-3.5 Turbo over GPT-4 in the Custom dataset requires cautious interpretation. Further analysis is needed to determine if this is due to actual model advancements or the limited sample size.
4.4. [RQ4]: Bug Detection.
[Experiment Goal]: Our aim is to evaluate to what extent the test suites generated by LLMs can detect bugs compared with those produced by an SBST method.
[Experiment Design]:
We employed two controlled experiments to assess the bug detection capabilities of the targeted models. The first experiment leveraged the Defects4J benchmark (Just et al., 2014), a well-established dataset encompassing 403 real-world software bugs across diverse Java projects. This experiment compared the performance of three models: GPT-3.5 turbo, Mistral 7B, and EvoSuite, a leading SBST tool. We rely only on Defects4J since SF110 does not provide ground truth fixed versions of programs to validate bug detection.
The second experiment builds on a Custom dataset we developed, containing 20 real-world software bugs extracted from recent (post-May 2023) GitHub projects. This dataset was necessary to evaluate the generalizability of the findings on unfamiliar datasets for the LLMs. The second experiment included four models: GPT-3.5 turbo, GPT-4 (a more advanced version of GPT-3.5), Mistral 7B, and Mixtral 8x7B (a potentially higher-capacity variant of Mistral 7B that we use sparingly due to budget constraints).
The evaluation procedure involved running the test suite on the buggy version of the class. If the test suite passes, the bug is undetected; if it fails, it is run on the fixed version. A passing result on the fixed version confirms bug detection. If it still fails, error messages and the number of failing tests are analyzed. A decrease in failing tests on the fixed class indicates successful bug detection; otherwise, the bug remains undetected.
[Experiment Results]: Table 8 presents the results of bug detection with the test suites generated by Mistral 7B, GPT 3.5-turbo and EvoSuite on the Defects4J dataset. GPT 3.5-turbo and Mistral 7B showed varied effectiveness across different prompting techniques, with CoT achieving the highest bug detection rate. Specifically, CoT with Mistral 7B detected 150 out of 403 bugs, outperforming other techniques. In contrast, EvoSuite detected only 3 out of 403 bugs. This discrepancy raises questions about the conventional applications and configurations of SBST tools like EvoSuite, which are primarily designed for regression testing scenarios. These tools may not be optimized for detecting bugs in a forward testing scenario, as evidenced by the low detection rates observed.
On the Custom dataset, all LLMs failed to detect bugs. This failure suggests that either the Defects4J dataset, in contrast to CMD, includes easy-to-detect bugs or the test cases of Defects4J have been memorized by the LLMs. LLMs also likely falter when prompted on novel software frameworks or languages that introduce less common syntax or constructs. Despite many classes in the custom dataset being of low complexity (Table 2), moderately and highly complex classes suggest that inherent complexities might further challenge effective bug detection. Finally, the high bug detection rates with ZSL prompts, contrary to FSL and even ToT, also cast suspicion of data leakage. CoT prompts exhibit the strongest overall performance regarding bug detection capabilities of the generated test suites. This is likely due to CoT’s effective simulation of real-world testing scenarios, which encourages deeper exploration of code logical flows to uncover bugs.
| Projects | # bugs | Mistral 7B | GPT 3.5-turbo | EvoSuite | ||||||||
| CoT | ToT | GToT | FSL | ZSL | CoT | ToT | GToT | FSL | ZSL | SBST | ||
| Cli | 30 | 17 | 14 | 14 | 14 | 16 | 20 | 19 | 14 | 18 | 20 | 1 |
| Closure | 63 | 22 | 11 | 10 | 10 | 17 | 18 | 8 | 11 | 5 | 12 | 0 |
| Codec | 8 | 3 | 2 | 2 | 2 | 2 | 3 | 2 | 2 | 2 | 2 | 0 |
| Collections | 1 | 1 | 1 | 1 | 1 | 1 | 1 | 1 | 1 | 1 | 1 | 0 |
| Compress | 22 | 10 | 9 | 9 | 8 | 10 | 10 | 10 | 11 | 9 | 12 | 1 |
| Csv | 9 | 5 | 5 | 5 | 5 | 6 | 6 | 4 | 5 | 4 | 6 | 0 |
| Gson | 9 | 4 | 0 | 0 | 0 | 3 | 4 | 0 | 0 | 0 | 2 | 0 |
| JacksonCore | 9 | 0 | 1 | 0 | 2 | 1 | 1 | 2 | 2 | 2 | 2 | 0 |
| JacksonDatabind | 68 | 1 | 0 | 0 | 1 | 1 | 0 | 0 | 0 | 0 | 1 | 0 |
| JacksonXml | 2 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
| Jsoup | 61 | 41 | 28 | 28 | 26 | 33 | 41 | 26 | 26 | 26 | 28 | 0 |
| JxPath | 14 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
| Lang | 14 | 12 | 11 | 11 | 11 | 12 | 11 | 10 | 10 | 10 | 11 | 0 |
| Math | 54 | 15 | 6 | 6 | 7 | 11 | 9 | 5 | 6 | 5 | 7 | 0 |
| Mockito | 33 | 19 | 5 | 5 | 6 | 9 | 19 | 5 | 4 | 6 | 6 | 0 |
| Time | 6 | 0 | 0 | 0 | 0 | 0 | 0 | 1 | 0 | 0 | 0 | 1 |
| Total | 403 | 150 | 93 | 91 | 93 | 128 | 143 | 93 | 92 | 88 | 110 | 03 |
1
Models include Mistral 7B, GPT 3.5-turbo, and EvoSuite. ZSL: Zero-shot learning, FSL: Few-shot learning, CoT: Chain-of-Thought, ToT: Tree-of-Thoughts, GToT: Guided Tree-of-Thoughts, SBST: Search-Based Software Testing.
5. Threats to Validity
This section addresses the threats to the validity of this study.
External Validity. The generalizability of our findings to real-world software testing scenarios might be limited by several factors. Our chosen benchmarks (Defects4J, SF110, and a custom dataset) may not fully capture the diverse complexities encountered in real-world software. Additionally, overestimation of LLM capabilities could occur if the models were trained on test cases included within the datasets. We mitigated this risk by creating a novel dataset, but further investigation is warranted. Focusing on four instruct-tuned LLMs from two families (OpenAI and MistralAI) restricts the representativeness of our results to the broader landscape of available LLMs. Furthermore, limitations exist regarding knowledge cutoff dates for the Mistral models. While the official paper lacks details, external sources suggest a 2021-2022 cutoff for Mistral 7B (Otterly, 2024; Saket3199, 2024). For Mixtral 8x7B, we only have the release date (December 2023), which may not reflect its actual knowledge cutoff. It’s crucial to acknowledge that LLMs are constantly evolving. Our findings represent their capabilities only during the study period (November 2023 to April 2024).
Internal Validity. Threats to internal validity in our study stem from the randomness of LLMs and EvoSuite, prompt design, and our bug detection methodology. To address randomness, we conducted each experiment thirty times and employed detailed statistical analysis to ensure reliable interpretations. Our custom-designed prompts, while based on model guidelines and previous research, might not be perfectly optimized for each model or test generation task. Our bug detection methodology mainly focuses on functional correctness, which may miss performance or security-related bugs. We also need to consider flaky tests, as they can result in misleading bug detection outcomes.
6. Related work
Siddiq et al. (Siddiq et al., 2023) evaluated the ability of LLMs like Codex, GPT-3.5-Turbo, and StarCoder to generate JUnit tests for Java, using benchmarks from HumanEval and EvoSuite SF110. They used zero-shot prompts and focused on aspects like compilation rates, correctness, coverage, and test smells. Our research aligns with their findings, highlighting similar challenges faced by LLMs. However, we further explored prompt engineering techniques to enhance test quality and coverage. Additionally, our study investigated bug detection capabilities, offering a more comprehensive view of LLMs’ applications in software testing.
Tang et al. (Tang et al., 2024) recently evaluated the effectiveness of ChatGPT and EvoSuite in generating Java unit test suites, focusing on correctness, readability, code coverage, and bug detection. Similarly to us, they find ChatGPT underperforms compared to EvoSuite in test suite effectiveness, especially in terms of code coverage. Surprisingly, they have claimed that EvoSuite test suites finds more bugs than LLM-generated test suites. This finding contradicts with our study (on slightly different projects’ code), potentially due to the methodology for bug detection: several works in the literature conclude on bug detection as soon as a test execution fails. Our definition is more constrained. Our study further goes beyond ChatGPT and EvoSuite, assesses various prompt designs, and evaluates on a comprehensive set of testing metrics.
Taking into account the limitations of existing LLMs, Rao et al. (Rao et al., 2023) have proposed to train a specific GPT-style language model, CAT-LM, for the task of test generation. To that end they align code and test files during training. With 2.7 billion parameters, CAT-LM excels in generating syntactically valid tests that closely mimic developer-written tests in structure and style. Unfortunately, CAT-LM has not explored the potential of prompt design in improving the capabilities of LLMs and has not targeted bug detection as a performance criterion.
7. Conclusion
We conducted a large-scale study investigating LLMs for unit test generation. Our experimental results record for the first time in the literature on applying LLMs for test generation to what extent prompt design significantly impacts test quality: chain-of-thought prompting techniques and our newly designed Guided Tree-of-though prompting strategy yield the most effective tests. While correctness issues persist, compilable LLM-generated tests resembled human-written ones in terms of adherence to coding styles, suggesting potential. A trade-off emerged: EvoSuite achieved higher coverage, while CoT-prompted GPT-3.5 Turbo and Mistral 7B excelled in bug detection. However, this success was primarily observed on known datasets, hinting at some limitations of LLMs with unfamiliar code. These promising yet preliminary results call for further research by the software engineering community to unlock LLMs’ full potential, particularly in generating tests that are relevant to the ultimate task of bug detection.
Open Science To promote transparency and facilitate reproducibility, we make our artifacts available to the community at:
https://anonymous.4open.science/r/LLM4TS-0F76/
The repository includes all LLM and SBST-generated tests, our Custome dataset, and all scripts for the replication of our study.