### v2

# Large-scale, Independent and Comprehensive study of the power of LLMs for test case generation

Abstract.
Unit testing, crucial for ensuring the reliability of code modules, such as classes and methods, is often overlooked by developers due to time constraints. Automated test generation techniques have emerged to address this, but they frequently lack readability and require significant developer intervention. Large Language Models (LLMs), such as GPT and Mistral, have shown promise in software engineering tasks, including test generation, but their overall effectiveness remains unclear.
This study presents an extensive investigation of LLMs, evaluating the effectiveness of four models and five prompt engineering techniques for unit test generation. We analyze 216 300 tests generated by the selected advanced instruct-tuned LLMs for 690 Java classes collected from diverse datasets. Our evaluation considers correctness, understandability, coverage, and test smell detection in the generated tests, comparing them to a widely used automated testing tool, EvoSuite. While LLMs demonstrate potential, improvements in test quality—particularly in reducing common test smells—are necessary. This study highlights the strengths and limitations of LLM-generated tests compared to traditional methods, paving the way for further research on LLMs in test automation.
1. Introduction
Unit tests plays a crucial role in software development, as they validate individual functions or code units, ensuring their correctness and reliability (Beck, 2000; Siddiqui, 2021; Shore and Warden, 2021). In particular, unit tests help in early bug detection, enhance code quality, provide documentation, support debugging, and reduce the risk of costly errors in production. However, manual unit test creation can be arduous and discouraging, leading some developers to skip testing entirely (Daka and Fraser, 2014; Kochhar et al., 2013b, a).
To cope with the challenges of test creation by developers, automated test generation techniques have gained traction, including methods like fuzzing (Miller et al., 1990; Zeller et al., 2019), feedback-directed random test generation (Toffola et al., 2017; Pacheco and Ernst, 2007; Pacheco et al., 2008), dynamic symbolic execution (Godefroid et al., 2005; Sen et al., 2005; Cadar et al., 2008), and search-based and evolutionary approaches (Fraser and Arcuri, 2011a, b). These techniques aim to create tests that maximize coverage through static or dynamic analyses. EvoSuite (Fraser and Arcuri, 2011b), a prominent search-based method, uses evolutionary algorithms to generate test suites for Java classes, aiming to maximize coverage. However, despite its effectiveness, the generated tests often lack readability (Almasi et al., 2017; Grano et al., 2018; Roy et al., 2020), contain obscure variable names (Daka et al., 2015), lack assertions (Panichella et al., 2020) or include overly general or excessive assertions (Panichella et al., 2020), and have shortcomings in detecting real bugs (Bacchelli et al., 2008; Pinto and Vergilio, 2010; Shamshiri et al., 2015; Almasi et al., 2017). Although these tests can inspire manual development of high-coverage test suites, they often lack naturalness and practical usability (Yuan et al., 2023; Shin et al., 2023; Almasi et al., 2017; Panichella et al., 2020; Alagarsamy et al., 2023).
Given the limitations of traditional analysis-based test generation methods, recent research has focused on using advanced deep learning techniques to improve unit test generation. Some techniques treat test generation as a neural machine translation problem, fine-tuning models to translate classes or methods into corresponding test prefixes or assertions (Tufano et al., 2020; Watson et al., 2020; Dinella et al., 2022; Nie et al., 2023; Shin et al., 2023; Alagarsamy et al., 2023). Recently, Large Language Models (LLMs), such as GPT-3 (Brown et al., 2020), have been leveraged for this purpose due to their extensive parameter sets and versatility (Fan et al., 2023). Studies (Wang et al., 2024; Fan et al., 2023; Kang et al., 2023; Lemieux et al., 2023; Chen et al., 2023; Siddiq et al., 2023) have then shown that LLMs can significantly enhance the quality of automated test case generation, streamlining the process for developers. These LLM-based test generation approaches are claimed to address issues of readability, understandability, and meaningfulness (Siddiq et al., 2024b; Yuan et al., 2023; Tang et al., 2024). The emergence of LLMs therefore may mark a significant evolution in software development, enabling more automated, intelligent, and efficient testing processes.
Despite the potential of advanced LLMs in software engineering tasks, it is crucial to critically evaluate their actual effectiveness in generating unit test cases. We are concerned in several main dimensions:
Data leakage: Sallou et al. (Sallou et al., 2023) caution that current methods might miss critical information and use potentially irrelevant metrics, with data leakage risks leading to overly optimistic results. There is thus a need to assess the capability of LLMs with ”unseen” code for test generation.
LLM diversity: The promise of LLMs for test generation has been mostly studied based on the capabilities of GPT-3, which is a closed model. Other models, such as Mistral 7B (Jiang et al., 2023), may be relevant to assess for this daily task in software engineering.
Prompting strategy: A recent survey by Wang et al. (Wang et al., 2024) has revealed that Zero-shot and Few-shot learning are the two most applied prompting techniques for LLM-based test case generation. Yet, other prompting strategies (Amatriain, 2024), such as Chain-of-Thought (Wei et al., 2022) and Tree-of-Thoughts (Yao et al., 2024), have already shown to substantially improve results in various software engineering tasks.
Evaluation methodology: Beyond classical metrics such as code coverage, few works actually attempt to assess whether the LLM-generated test cases are readable or contribute to the presence of test smells. In our study, we go beyond these traditional metrics by focusing on the prevalence of test smells, which are indicative of poor design and maintainability in test suites.
This paper: Our work investigates the effectiveness of LLMs in unit test generation and how prompt engineering techniques might affect their performance. Drawing inspiration from (1) works that compare Search-Based Software Testing (SBST) against ChatGPT (Tang et al., 2024; Siddiq et al., 2024b), (2) comprehensive surveys in the field (Wang et al., 2024), and (3) insightful early experiments on the threats of using LLMs in software engineering (Sallou et al., 2023), we undertake a large-scale study to evaluate the potential of LLMs and prompt engineering to improve unit test case generation.
Unlike much of the prior research, which focuses mainly on criteria related to syntax accuracy, compilability, and at best, code coverage, our study further evaluates these criteria along with the capability of LLMs to generate test suites that are free from common test smells.
We assess four popular LLMs, GPT (3.5 and 4) from OpenAI and Mistral (7b and 8X7B) from MistralAI, and employ five prompt techniques implemented in a way that is suitable for test generation. Our comparative analysis examines the unit test suites generated by these LLMs against those from EvoSuite, a widely-known SBST tool. We rely for our experiments on two well-known datasets, SF110 (Fraser and Arcuri, 2014) and Defects4J (Just et al., 2014), as well as a new dataset, collected based on the criterion111This Custom dataset includes projects released on GitHub after May 2023, to ensure the study’s findings are pertinent and free from biases associated with pre-existing LLM training data. that the test cases must have been “unseen” by the LLMs during their training.
Our main contributions are as follows:
We provide insights on the capabilities of LLMs for unit test generation based on a comprehensive analysis of the impact of prompt engineering. To the best of our knowledge, this constitutes the first work in the literature that addresses this question.
We describe our findings based on a comprehensive array of evaluation metrics, including correctness, readability, code coverage, and the prevalence of test smells. By leveraging various datasets and models, we provide a robust basis for comparisons and future research. In particular, our assessment on an ”unseen” dataset strengthens our study’s relevance and mitigates biases related to data leakage in LLMs.
We provide a comparative evaluation of LLMs against a traditional SBST method.
To aid ongoing and future research in this field, we release a comprehensive benchmark including (1) a labeled dataset of 216,300 tests generated by four instruction-tuned LLMs across 690 Java classes, (2) and performance evaluation scripts. These artifacts are available at:
The remainder of this paper is structured as follows: Section 2 details the background of LLMs for unit test generation, and prompt engineering. Section 3 outlines our experimental setup, followed by Section 4, which discusses the evaluation results. While Section 5 addresses potential study threats and limitations, Section 6 reviews related work. Section 7 discusses the findings, their implications, and takeaways. Section 8 offers concluding remarks.
2. Background
In this section, we introduce the fundamental concepts essential for understanding our paper. We succinctly cover LLMs for unit test generation along with prompt engineering, which constitutes the central ideas of this study.
2.1. LLM-based Unit Test Generation
Large language models (LLMs) are extensive neural networks built primarily on the Transformer architecture (Vaswani et al., 2017) and trained on vast text corpora and code using self-supervised techniques like Masked Language Modeling and Causal Language Modeling (Shanahan, 2024; Naveed et al., 2023; Hadi et al., 2023). These models, including encoder-only like BERT (Devlin et al., 2018), decoder-only like GPT (Brown et al., 2020; Yenduri et al., 2024), and encoder-decoder models like T5 (Raffel et al., 2020), are often enhanced with methods like Reinforcement Learning from Human Feedback (RLHF) (Ziegler et al., 2019; Bai et al., 2022) to better align with human intentions. When interacting with these models, the instructions provided alongside any input data form what is referred to as a prompt (Siddiq et al., 2024b; Amatriain, 2024; Vogelsang and Fischbach, 2024; Siddiq et al., 2024a). Indeed, a prompt is a text description in natural language that specifies the task to be performed by an LLM. As shown in Figure 1, an LLM can be prompted 222https://chatgpt.com/share/99d6468e-98ed-40b1-9b14-17bb9c239f9c to generate a test suite for a Divider Class, and it will respond by producing the test suite, illustrating its ability to understand and execute the task specified in the prompt.
LLMs have been notably applied in various fields, with several specialized models tailored for software engineering tasks like code synthesis and test generation (Fan et al., 2023; Wang et al., 2024). Prominent examples, such as Codex (Chen et al., 2021) and StarCoder (Li et al., 2023), can perform test generation. Despite the widespread usage of LLMs in code and test generation tasks (Wang et al., 2024; Fan et al., 2023; Yuan et al., 2023; Siddiq et al., 2024b; Schäfer et al., 2023; Chen et al., 2023; Tang et al., 2024), their effectiveness in unit test generation through prompt engineering remains under-explored. This study addresses this gap by evaluating the capabilities of LLMs in generating unit tests, marking a significant exploration into their application in software engineering.
2.2. Prompt Engineering
Prompt engineering involves creating prompts which enable LLMs to effectively address specific tasks (Amatriain, 2024; Siddiq et al., 2024a; Vogelsang and Fischbach, 2024; Naveed et al., 2023; Sahoo et al., 2024). This technique is particularly useful in complex problem-solving scenarios, where a single line of reasoning might not be enough (Amatriain, 2024; Kojima et al., 2022; Wei et al., 2022). It allows LLMs to adopt a more human-like approach by considering multiple possibilities before reaching a conclusion. This emerging field involves crafting prompts that guide LLMs to generate desired outputs, ranging from broad descriptions to detailed specifications. The importance of prompt engineering is twofold: Firstly, highly specific prompts yield precise responses, while more general prompts may produce less relevant outputs (Si et al., 2022). Secondly, adjusting prompt instructions can help control the output format of LLMs, minimizing the need for post-processing and improving performance evaluation (Macedo et al., 2024).
Various studies (Fan et al., 2023; Wang et al., 2024) have leveraged LLMs and prompt engineering for automated software engineering tasks, commonly using basic approaches such as Zero-shot (Kojima et al., 2022) and Few-shot learning (Brown et al., 2020). Recently, research has broadened prompt engineering to include techniques such as Chain-of-Thought (Wei et al., 2022), self-consistency (Wang et al., 2022), and Tree-of-Thoughts (Wang et al., 2022). Figure 2 provides examples of the prompts used for test generation, which were carefully crafted following prompt engineering best practices and insights from prior research (Tang et al., 2024; Chen et al., 2023; Tufano et al., 2020) (additional examples of prompt engineering techniques are available in the replication package).
Zero-shot learning enables models to generalize to new tasks without specific examples but may falter with complex tasks (Sahoo et al., 2024; Kojima et al., 2022). Few-shot learning, or in-context learning, uses a few examples to improve model performance on novel tasks (Brown et al., 2020; Reynolds and McDonell, 2021). Chain-of-Thought (CoT), introduced by Wei et al. (Wei et al., 2022), outlines intermediate reasoning steps and can be enhanced by combining it with few-shot prompts for complex reasoning. Self-consistency, proposed by Wang et al. (Wang et al., 2022), uses diverse reasoning paths to select the most consistent answers, improving deep reasoning tasks. Tree-of-Thoughts (ToT), a newer technique proposed by Yao et al. (Yao et al., 2024) and (Long, 2023), builds on Chain-of-Thought by employing a systematic exploration through a ”tree” of reasoning steps, integrating search algorithms for more strategic problem-solving. This study explores various prompt engineering methods to adapt pre-trained LLMs for unit test generation, emphasizing their application in guiding LLM behavior in test generation tasks.
2.3. Test Smells
The primary goal of unit testing is to ensure that individual components of a program function as expected and meet specified requirements (Koomen and Pol, 1999). A unit typically refers to an isolated segment of code, such as a method or class, that can be tested independently. Like production code, unit tests need to be not only correct but also maintainable and readable (Gonzalez et al., 2017). Test smells are indicators of potential issues or poor practices within a unit test suite (Greiler et al., 2013; Guerra and Fernandes, 2007; Palomba et al., 2016; Peruma et al., 2019). These smells may signal inefficiencies, complexity, or excessive coupling to the underlying code, making the tests harder to maintain or understand (Meszaros et al., 2003). Examples include tests that are overly fragile, difficult to understand, or tightly bound to implementation details. For instance, the Java code in Listing 1 and Listing 2 from the TsDetect documentation 333https://testsmells.org/pages/testsmellexamples.html highlights two common test smells: Assertion Roulette and Magic Number Test. These test smells frequently appear in automatically generated tests, such as those produced by tools like EvoSuite or LLM-based test generation.
Assertion Roulette occurs when multiple assertions are made in a test without clear explanations, making it difficult to identify which assertion caused the test to fail. In Listing 1, the method assertThat() is called three times, each testing a different condition, but without any explanation messages. If one of these assertions fails, identifying the source of the failure becomes challenging. Magic Number Test refers to the use of unexplained numeric literals in test assertions, making it unclear what the values represent. In Listing 2, the test uses numeric literals in the assertions (15 and 30) without explaining their significance. While the test may be syntactically correct and compilable, the lack of clarity hinders maintainability and understandability.
Although these tests are functionally correct, their test smells make them harder to maintain and debug, presenting challenges in long-term software development.
3. Experimental Setup
3.1. Overview
Our methodology, illustrated in Figure 3, involves prompting LLMs with diverse datasets to generate unit test cases, followed by evaluation for syntax correctness, compilability, and validation. Any substandard outputs are discarded, and the process is repeated thirty times to ensure reliability. To ensure a fair comparison, we adopted the same approach as EvoSuite, which was executed thirty times for each class in (Tang et al., 2024).
Our study focuses on generating unit tests at the class level, reflecting real-world testing scenarios where testing an entire class provides a more comprehensive evaluation of the model’s capabilities. Class-level testing captures the interactions between methods within a class and presents more complex challenges than method-level testing. By focusing on the class level, we aim to assess how well LLMs can generate tests that account for dependencies and interactions within a class. We investigate the quality of LLM-generated tests in terms of output format, syntax correctness, and compilability (RQ1), as well as comprehensibility (RQ2). We also compare them with EvoSuite in terms of code coverage (RQ3) and test smell detection (RQ4).
3.2. Research Questions
This study evaluates the effectiveness of LLMs in test case generation and examines the impact of prompt engineering on their performance. It includes an investigation structured around four specific Research Questions (RQs).
RQ1: How does prompt engineering influence the generation of syntactically correct and compilable test suites by LLMs? We investigate the potential for suitable prompt engineering to guide LLMs towards producing test suites that: ❶ are syntactically correct: this ensures the code adheres to the programming language’s rules, increasing the likelihood of successful compilation; and ❷ compile without errors: compiled code enables execution but doesn’t guarantee functionality. Findings in the study of this research will contribute to improving the ability of LLMs to generate well-formed test suites that can be automatically extracted for code analysis, ultimately promoting development efficiency.
RQ2: Beyond syntactic correctness, how well do LLMs generate human-readable and maintainable test suites? This research question examines the readability and maintainability of test suites produced by LLMs. Well-written test suites are essential for developers to: ❶ Understand the test logic: Clear and concise code makes it easier to grasp the purpose and functionality of the tests; ❷ Maintain the tests over time: As code evolves, readable tests require less effort to update and keep current; and ❸ Integrate tests into existing codebases: Easy-to-understand tests facilitate smooth merging with existing code. By evaluating these aspects in this research question, we aim to determine whether LLM-generated test suites are practical for real-world development workflows.
RQ3: How does the code coverage achieved by LLM-generated test suites compare to those generated by a Search-Based Software Testing technique? This research question investigates the effectiveness of LLMs in generating test suites that thoroughly cover the code under test. Code coverage is a key metric indicating the percentage of code executed by the test suite. Higher coverage suggests a more comprehensive evaluation, potentially leading to better bug detection. EvoSuite, a prominent SBST tool, serves as a benchmark to assess the capabilities of LLMs in this domain.
RQ4: How do LLM-generated test suites compare to EvoSuite-generated test suites in terms of test smell prevalence? This research question explores the quality of LLM-generated test suites by analyzing the presence of test smells—indicators of poor design and maintainability. We compare the absence of test smells of LLM-generated test suites against those produced by EvoSuite. The goal is to determine whether LLMs, through prompt engineering, can produce tests that are less prone to common test smells and thus more maintainable than those generated by traditional SBST techniques like EvoSuite.
3.3. Large Language Models
Our study focuses on four instruction-tuned decoder-only models known for their enhanced ability to follow instructions compared to base models. These models belong to the OpenAI and MistralAI families and are detailed in Table 1.
One such model, GPT-3.5-turbo, has a token limit of 4,096. Research suggests that exceeding this limit when splitting a class into test case generation methods can hinder the model’s understanding of the class context, potentially reducing test quality (Tang et al., 2024). To ensure a fair comparison, all models in our study are restricted to this token limit, and any inputs exceeding it are excluded. Token counts are determined using OpenAI’s tiktoken444https://github.com/openai/tiktoken library.
Furthermore, all models are set to a standard temperature of 0.7. Our selection criteria prioritized models that excel in code generation and possess accessible APIs, making them suitable for both academic research and industrial applications (Wang et al., 2024; Fan et al., 2023).
| Dataset | Description | Number of Projects | Number of Classes | Number of Bugs | Complexity Overview | ||
| Low | Moderate | High | |||||
| SF110 | Derived from DynaMOSA benchmark (Panichella et al., 2017), includes 346 Java classes from 117 projects. | 74 | 182 | N/A | 170 | 138 | 120 |
| Defects4J | Contains 835 reproducible bugs across 17 open-source projects, focuses on detailed test suites to expose bugs. | 17 | 477 | 403 | 254 | 177 | 130 |
| Custom mini Dataset (CMD) | Includes ODC a collaborative project to enhance database functionalities and Conductor OSS an open-source project management tool aimed at efficiency and scalability. | 2 | 31 | 20 | 16 | 9 | 5 |
3.4. Prompt Design
Unit test generation research highlights the importance of practical and focused methods that avoid overly complicated solutions (e.g., (Siddiq et al., 2024b; Yuan et al., 2023; Chen et al., 2023; Tang et al., 2024)). To achieve this, we designed prompts with varying complexity levels based on a clear and well-structured template. Each prompt consists of two parts: a Natural Language Description (NLD) outlining the task and a Source Code placeholder to insert the Class Under Test (CUT). We will now explore the details of each part, explaining how they align with established unit testing practices.
Natural Language Description: Drawing on established best practices for prompt effectiveness, we designed our prompts as Natural Language Descriptions (NLDs) that outline the task for LLMs. These NLDs incorporate three key elements: (i) a role-playing directive to mimic a professional software tester (or several testers), which is known to improve test generation (Tang et al., 2024; Chen et al., 2023); (ii) a task-specific directive to generate a complete Junit 4 test file for a specified class, and (iii) a control statement requiring the output to be enclosed between specific start and end markers, ensuring structured output.
Source Code: Consistent with existing studies on learning-based unit test generation (Tufano et al., 2020), our prompts vary based on the engineering technique, either including (i) just the class name and its source code (Figure 2(a)), or (ii) the class name, its source code, plus an example of another class’s source code and its test suite, to provide models with context or comparative examples (Figure 2(b)).
This study investigates the impact of varying prompt complexity on the quality of generated test cases. We employ five distinct prompt engineering techniques: Zero-shot Learning (ZSL), Few-shot learning (FSL), Chain-of-Thought (CoT), Tree-of-thoughts (ToT), and a novel technique called Guided Tree-of-thoughts (GToT). ZSL provides prompts without specific examples, forcing models to generalize to new tasks. FSL offers a few examples to guide the model towards better performance. For Few-shot learning, we chose a medium-sized human-written class and its test suite to fit within the LLM token limit. This pair was used consistently across all experiments. CoT structures prompts to encourage logical reasoning, while ToT promotes exploration of multiple reasoning paths for more informed decisions. GToT attempts to exploit advantages of both CoT and ToT by guiding the reasoning of each solicited expert within the prompt itself. The main intuition is that combining the strengths of CoT and ToT in a single prompt could improve LLMs performances. While CoT instructs a single tester to list methods, generate basic and edge case tests, and compile them into a JUnit file, GToT simulates three testers collaborating to refine test cases at each step, resulting in more comprehensive and accurate tests. This study focuses on these techniques as test case generation demands a deep understanding of the code and the ability to reason through complex scenarios. Further details and comparisons of these prompt templates are available in the replication package.
3.5. The EvoSuite SBST tool
EvoSuite Configuration
This section details the configuration used for EvoSuite, a well-established Search-Based Software Testing (SBST) tool, in our study. We employed EvoSuite with its default settings, including the DynaMOSA algorithm (Panichella et al., 2017). This algorithm focuses on achieving high coverage for branches, methods, and lines of code within the test suite.
To account for the inherent randomness of genetic algorithms used by EvoSuite, we executed the tool 30 times for each class under test like in (Tang et al., 2024). We allocated a three-minute time budget for each test case generation, chosen to reflect practical constraints typically encountered by developers. Previous studies (Fraser et al., 2015; Arcuri and Fraser, 2013; Shamshiri et al., 2015) have shown that a three-minute time budget is generally sufficient for effective test case generation.
Hardware
For reference, EvoSuite experiments were conducted on an AMD EPYC 7552 48-Core Processor running at 3.2 GHz with 640 GB of RAM. However, it’s important to note that not all available resources were necessarily utilized during the tests.
3.6. Datasets
To minimize selection bias and ensure generalizability, we employed three diverse datasets: SF110, Defects4J, and a Custom Mini Dataset (CMD). Both SF110 and Defects4J were used to evaluate both EvoSuite and LLM performance. The CMD, containing GitHub projects created after May 2023, was exclusively tested with LLMs due to compatibility issues with EvoSuite’s Java version requirements (see Issue #433 on the EvoSuite GitHub repository(evo, [n. d.])). This dataset focused on well-maintained projects from the OceanBase Developer Center (ODC) (OceanBase, 2023) and Conductor OSS (conductor oss, 2023).
ODC, with 13 releases, 527 stars, and 20 contributors, is a collaborative database development tool offering features such as SQL development, risk management, and data security. Conductor OSS, originally developed by Netflix, is a microservices orchestration platform with 11 releases, 15.5k stars, and 53 contributors. These projects were selected due to their complexity and active development, providing a robust basis for testing LLM-generated unit tests. CMD was limited to two projects to accommodate the token limits of the LLMs, but future work will expand the dataset to improve generalizability.
The datasets encompass projects with code of varying levels of cyclomatic complexity (detailed in Table 2), ranging from low to high. This diversity helps mitigate potential overfitting on specific complexity profiles and provides valuable insights into the potential of LLMs for real-world automated test case generation.
Notice. Due to budgetary reasons, we only run GPT-4 and Mixtral-8x7B on the smallest dataset (CMD). Other models (GPT-3.5-Turbo and Mistral-7B) are run on all datasets.
4. Experimental Results
| Model Name | # of Generated Files | # of Files with code | MSR | CSR | ||||||||
| SF110 | Defects4J | CMD | SF110 | Defects4J | CMD | SF110 | Defects4J | CMD | SF110 | Defects4J | CMD | |
| GPT 3.5-turbo | 27 300 | 71 550 | 4 650 | 23 048 | 60 565 | 4 633 | 84.42% | 84.65% | 99.64% | 78.10% | 78.34% | 90.79% |
| GPT 4∗1 | – | – | 4 650 | – | – | 4 356 | – | – | 93.68% | – | – | 83.59% |
| Mistral 7B | 27 300 | 71 550 | 4 650 | 25 664 | 65 180 | 4 226 | 94.00% | 91.10% | 90.88% | 79.97% | 80.35% | 83.61% |
| Mixtral 8x7B∗1 | – | – | 4 650 | – | – | 2 947 | – | – | 63.38% | – | – | 53.57% |
| Total/Average (%) | 54 600 | 143 100 | 18 600 | 48 712 | 125 745 | 16 162 | 89.21 | 87.88 | 86.90 | 79.04 | 79.35 | 77.89 |
1
∗Note that due to budget constraints, we only paid to query GPT-4 and Mixtral 8x7B for CMD.
| Model Name | Syntactically Correct/Compilable Code SF110 (%) | Syntactically Correct/Compilable Code Defects4J (%) | Syntactically Correct/Compilable Code CMD (%) | ||||||||||||
| ZSL | FSL | CoT | ToT | GToT | ZSL | FSL | CoT | ToT | GToT | ZSL | FSL | CoT | ToT | GToT | |
| GPT 3.5-turbo | 21.30/5.94 | 17.30/4.21 | 18.05/5.46 | 19.70/4.08 | 23.65/9.67 | 21.51/2.95 | 16.43/2.22 | 17.51/3.23 | 20.02/2.26 | 24.54/6.76 | 21.70/2.70 | 20.77/2.44 | 20.09/5.92 | 16.39/2.68 | 21.06/3.46 |
| GPT 4 | –/– | –/– | –/– | –/– | –/– | –/– | –/– | –/– | –/– | –/– | 23.00/1.78 | 22.90/2.21 | 20.68/2.65 | 17.60/1.90 | 15.82/2.39 |
| Mistral 7B | 21.25/1.09 | 21.15/1.04 | 20.95/0.91 | 20.43/1.74 | 16.23/3.37 | 20.26/0.45 | 20.23/0.43 | 20.65/0.27 | 20.13/0.44 | 18.73/1.09 | 20.37/0.00 | 20.29/0.03 | 20.60/0.00 | 20.11/0.03 | 18.62/0.31 |
| Mixtral 8x7B | –/– | –/– | –/– | –/– | –/– | –/– | –/– | –/– | –/– | –/– | 18.85/0.16 | 23.36/0.12 | 18.51/0.08 | 15.58/0.44 | 23.60/0.16 |
| Total/Average (%) | 21.27/3.52 | 19.22/2.62 | 19.50/3.19 | 20.06/2.91 | 19.94/6.52 | 20.89/1.70 | 18.33/1.32 | 19.08/1.75 | 20.08/1.35 | 21.63/3.93 | 21/1.16 | 21.83/1.20 | 19.97/2.16 | 17.42/1.26 | 19.78/1.58 |
1
∗ZSL: Zero-shot learning, FSL: Few-shot learning, CoT: Chain-of-Thought, ToT: Tree-of-Thoughts, GToT: Guided Tree-of-thoughts.
4.1. [RQ1]: Output Format, Syntactical Correctness, and Compilability.
[Experiment Goal]: We aim to explore how different prompts may lead the LLMs to exhibit difference effectiveness in producing syntactically correct and compilable test suites in the required output format.
[Experiment Design]: We consider five prompt designs to instruct the LLMs to generate test cases and use regular expressions to extract the source code of test suites from the LLM outputs. To evaluate the effectiveness of the techniques, we adopted the Match Success Rate (MSR) and Code Extraction Success Rate (CSR) metrics introduced by Macedo et al. (Macedo et al., 2024). MSR measures the proportion of outputs where the output could be matched by a pre-defined regexp pattern. It is thus a proxy to measure to what extent the LLM adheres to the format specified in the prompt. CSR computes the proportion of cases where a source code can be extracted within the matched output. We opted to use regular expressions as a lightweight and efficient solution. This approach is consistent with prior work (Macedo et al., 2024).
We analyze the extracted code using Javalang 0.13.0 (c2nes, 2012) to verify its syntactical correctness. If syntactically correct, the test suite then undergoes testing on the Java Virtual Machine (JVM) to check for compilability (for this RQ) as well as executability, and any runtime errors (in subsequent RQs). We also assess the quality of the generated test cases using SpotBugs 4.8.3 (spotbugs, 2016), an open-source static analyzer. SpotBugs checks Java programs for over 400 types of bugs and suboptimal coding practices. This analysis identifies how well the test cases detect common programming mistakes and ensure adherence to coding best practices.
[Experiment Results]: Our study generated a total of 216 300 files, of which 190 619 contained source code. Among these, 171 187 were syntactically correct test suites, while 25 681 files contained only textual outputs or incorrect data. Table 3 provides performance measurements of LLMs in terms of MSR and CSR.
GPT 3.5-turbo achieves the highest MSR (99.64%) and CSR (90.79%) on the Custom dataset. Mistral 7B also achieve high performance on the SF110 and Defects4J datasets. In contrast, Mixtral 8x7B struggles on the Custom dataset, in comparison to other models. When using GPT-4, we record a high CSR score at 83.59%.
Table 4 summarizes the performance of LLMs in generating syntactically correct (i.e., parsable) and compilable test suites. Overall, the metrics clearly show that LLMs underperform in generating syntactically correct test cases and are further struggling to generate compilable ones. In the best case scenario, GPT-3.5-Turbo with GToT prompting offers the highest effectiveness score on the Defects4J dataset with only 24.54% of generated test cases being syntactically correct, and 6.76% of these being also compilable. Overall, the results show that, on average, only 12% of the syntactically correct test suites compile successfully. The discrepancy between the total number of generated files and the number of syntactically correct test suites stems from the complexity of generating tests at the class level. Class-level generation, which involves generating tests for entire classes rather than individual methods, introduces additional challenges, including higher susceptibility to compilation errors. As a result, not all generated files contained valid or compilable test suites, contributing to the difference between the total generated files and the correctly compiled tests.
Following compilation, we perform static analysis using SpotBugs, which identifies 8 461 potential issues across 20 530 test cases (Table 5). We classify these issues (aka bugs) into four severity levels: Scariest (most critical, affecting functionality and security), Scary (significant issues if unaddressed), Troubling (minor issues), and Of Concern (indicating lower severity). The majority (55.58%) of the issues are labeled as only Of Concern, i.e., they are informational and of minimal risk. We find that Few-shot prompts lead the LLMs to generate most (61.72%) of the ’Scariest’ bugs. We nevertheless note that GToT is associated with most of the ’Scary’ (63.92%), ’Troubling’ (69.10%), and ’Of Concern’ (33.77%) bugs.
| Model | Dataset | Prompt | Classes Affected | |||
| Of Concern | Scariest | Scary | Troubling | |||
| GPT 3 | SF110 | ZSL | 600 | 0 | 0 | 0 |
| FSL | 410 | 410 | 0 | 0 | ||
| CoT | 553 | 0 | 0 | 0 | ||
| ToT | 20 | 0 | 1 | 0 | ||
| GToT | 825 | 0 | 825 | 825 | ||
| Defects4J | ZSL | 212 | 0 | 0 | 212 | |
| FSL | 177 | 177 | 0 | 0 | ||
| CoT | 219 | 0 | 0 | 0 | ||
| ToT | 21 | 1 | 1 | 2 | ||
| GToT | 242 | 0 | 0 | 0 | ||
| CMD | ZSL | 17 | 0 | 0 | 0 | |
| FSL | 18 | 0 | 0 | 0 | ||
| CoT | 22 | 0 | 0 | 0 | ||
| ToT | 3 | 0 | 0 | 0 | ||
| GToT | 20 | 0 | 0 | 0 | ||
| GPT 4 | CMD | ZSL | 15 | 0 | 0 | 15 |
| FSL | 14 | 0 | 0 | 14 | ||
| CoT | 0 | 0 | 0 | 0 | ||
| ToT | 7 | 0 | 0 | 1 | ||
| GToT | 19 | 0 | 0 | 0 | ||
| Mistral7B | SF110 | ZSL | 129 | 129 | 129 | 0 |
| FSL | 135 | 135 | 0 | 0 | ||
| CoT | 125 | 0 | 125 | 125 | ||
| ToT | 212 | 212 | 212 | 0 | ||
| GToT | 347 | 0 | 0 | 0 | ||
| Defects4J | ZSL | 48 | 48 | 0 | 0 | |
| FSL | 60 | 60 | 0 | 0 | ||
| CoT | 32 | 32 | 0 | 0 | ||
| ToT | 63 | 63 | 0 | 0 | ||
| GToT | 131 | 0 | 0 | 0 | ||
| CMD | ZSL | 0 | 0 | 0 | 0 | |
| FSL | 0 | 0 | 0 | 0 | ||
| CoT | 0 | 0 | 0 | 0 | ||
| ToT | 1 | 0 | 0 | 0 | ||
| GToT | 0 | 0 | 0 | 0 | ||
| Mixtral8x7B | CMD | ZSL | 2 | 0 | 0 | 0 |
| FSL | 0 | 0 | 0 | 0 | ||
| CoT | 0 | 0 | 0 | 0 | ||
| ToT | 0 | 0 | 0 | 0 | ||
| GToT | 4 | 0 | 0 | 0 | ||
1
∗ ZSL: Zero-shot Learning, FSL: Few-shot Learning, CoT: Chain-of-Thought, ToT: Tree-of-Thoughts, GToT: Guided Tree-of-Thoughts.
4.2. [RQ2]: Understandability.
[Experiment Goal]: We aim to assess the readability of code generated by LLMs through code style checking and understandability measures.
[Experiment Design]: We use Checkstyle 10.14.2 (checkstyle, 2001) to ensure compliance with Sun Code Conventions (Oracle, 1999) and Google Java Style (Google, 2015), verifying that LLM-generated test suites adhere to these standards. Additionally, we use metrics from Dantas et al. (Dantas and Maia, 2021) for evaluating cognitive and cyclomatic complexities with the PMD 7.0.0-rca tool (PMD, 2012). Cyclomatic complexity quantifies independent paths in the code, suggesting its complexity and error potential, while cognitive complexity focuses on the code’s ease of understanding, considering factors like structure and naming. These measures are essential for evaluating code maintainability and pinpointing potentially confusing or complex parts. The default thresholds in PMD for cyclomatic and cognitive complexity are 10 and 15, respectively.
Code Style Checking Results. Figure 4 visualizes the most common violations for each prompting technique according to both Sun and Google standards.
– Google Java Style: For Google coding standards, FSL achieved the best adherence with the fewest violations, followed by CoT. Other techniques showed lower adherence, reducing code readability. Common issues (Figure 4(a)) included incorrect indentation, improper import order, missing Javadoc documentation, wildcard imports, and long lines.
– Sun Code Conventions: For Sun coding standards, Few-shot learning is again the most compliant, followed by CoT and Zero-shot learning. ToT and GToT had higher violations, indicating weaker adherence and readability. Common issues (Figure 4(b)) included inadequate method documentation, use of ”magic” numbers, regex problems, excessive line length, and insufficient package documentation. These findings highlight the impact on code maintainability and clarity.
These evaluations help gauge the frequency and severity of style violations, with fewer average violations indicating better compliance with coding standards and improved readability. Detailed results for Checkstyle violations against Google and Sun Code Styles are documented in the replication package, along with distribution plots and statistical measures used to compare prompt strategies against coding standards.
Code Understanding Results. Table 6 shows the complexity Results by Prompt Engineering Techniques. Cyclomatic complexity helps assess the maintainability of the codebase, with higher values indicating lower maintainability, increased risk of errors, and greater difficulty in understanding and modifying the code. Lower cognitive complexity indicates more readable and understandable code. Note that PMD and Sonar, each with different levels of cyclomatic complexity. To facilitate comparison, we selected PMD levels but adjusted them according to Sonar’s standards (Campbell, 2023).
| Cyclomatic Complexity | ZSL | FSL | CoT | ToT | GToT |
| Low (1-4) | 3 601 | 2 794 | 3 669 | 2 960 | 7 468 |
| Moderate (5-7) | 4 | 11 | 6 | 6 | 1 |
| High (8-10) | 1 | 3 | 4 | 2 | 0 |
| Very High (11+) | 0 | 0 | 0 | 0 | 0 |
| Cognitive Complexity | ZSL | FSL | CoT | ToT | GToT |
| Low (¡5) | 3 600 | 2 804 | 3 674 | 2 967 | 7 469 |
| Moderate (5-10) | 4 | 4 | 5 | 1 | 0 |
| High (11-20) | 1 | 0 | 0 | 0 | 0 |
| Very High (21+) | 1 | 0 | 0 | 0 | 0 |
1
ZSL: Zero-shot learning, FSL: Few-shot learning, CoT: Chain-of-Thought, ToT: Tree-of-Thoughts, GToT: Guided Tree-of-Thoughts.
– Cognitive Complexity: Most test suites generated exhibit low complexity, particularly those produced by the Guided Tree-of-Thoughts (GToT) technique, which resulted in 7,469 low-complexity cases (Table 6(b)), the highest among all techniques. This accounts for over 99% of the tests generated by GToT, demonstrating its ability to consistently generate simple and maintainable code. Instances of moderate to very high complexity are rare, with only 1 case in the moderate range and none in the high or very high ranges, further highlighting the overall effectiveness of these techniques in producing straightforward and understandable code.
– Cyclomatic Complexity: Most generated test suites also exhibit low cyclomatic complexity, especially with the GToT technique, which produced 7,468 low-complexity cases (Table 6(a)), accounting for nearly all the test suites generated by this technique. Moderate and high complexity instances are minimal, with just 1 moderate-complexity case and no high-complexity cases. Other techniques like CoT and ZSL also result in a majority of low-complexity test suites, though they have slightly higher instances of moderate complexity (6 and 4 cases, respectively). This indicates that while most techniques generate maintainable code, GToT stands out for producing simpler, more maintainable test suites.
4.3. [RQ3]: Code Coverage.
[Experiment Goal]: We aim to evaluate how LLM-generated test suites fair against SBST-generated test suites in terms of code coverage.
[Experiment Design]: We consider EvoSuite as the baseline SBST tool for comparison. We measure line, instruction, and method coverage using the JaCoCo 0.8.8 tool (JaCoCo, 2022). For each class, we generate the test suites by each different LLM/prompt combination and Evosuite.
[Experiment Results]: We assess the effectiveness of each model/prompt combination or technique by employing statistical measures, such as sample size, minimum, maximum, quartiles, median, and mean, to compare code coverage metrics. Due to space constraints, all detailed results are provided in the replication package. Figure 5 illustrates example comparison results for Line coverage for Mistral-7B, GPT-3.5-Turbo and EvoSuite. We recall that since EvoSuite is incompatible with the Custom dataset (cf. Section 3.6), it was not applied for this dataset.
LLMs exhibit variable performance across different metrics, particularly GPT-3.5 Turbo, which is heavily influenced by advanced prompting techniques like GToT and CoT (Figure 5). In the Custom dataset, GPT-3.5 Turbo achieves near-perfect median coverage (100%), suggesting effective adaptation to novel software contexts. Interestingly, GPT-3.5 Turbo performs better than GPT-4 on this dataset, potentially due to the smaller size or specific limitations of GPT-4 for this particular task. However, GPT-3.5 Turbo’s performance varies across other datasets. It demonstrates a low median line coverage of 3.57% in SF110 but achieves significantly higher coverage in Defects4J (82.74% lines, 85.0% instructions, 100% methods) compared to SF110 (2.0% lines, 16.67% instructions). Mistral 7B also shows inconsistent performance, particularly struggling in SF110 with low median coverages (3.03% lines, 2.0% instructions, 20.0% methods). However, it achieves improved results in Defects4J, with higher median coverages (63.4% lines, 66.0% instructions, 100% methods).
Performance discrepancies between SF110 and Defects4J suggest the influence of dataset complexity or potential data leakage (discussed in Sallou et al. (Sallou et al., 2023)). Table 2 supports this notion, as SF110 exhibits the highest proportion of classes in moderate and high complexity categories despite its smaller sample size. Limited accessibility of SF110, unlike Defects4J’s broad availability (potentially inflating metrics as discussed in (Sallou et al., 2023)), might mitigate leakage risks. Favorable results in the new, smaller dataset could be skewed by its lower variability and simplicity. The dominance of low complexity classes in this dataset (Table 2) suggests potentially simpler and more modularized code, which may be easier to understand and less challenging for the models. EvoSuite consistently achieves near-perfect coverage across all metrics, demonstrating its effectiveness as a search-based software testing tool. In contrast, LLMs (e.g., GPT-3.5 Turbo and Mistral 7B) cover significant code portions but fall short of EvoSuite’s comprehensive levels. This stems from their differing test generation approaches: EvoSuite prioritizes maximizing coverage, while LLMs rely on learned patterns and instructions, which may not always guarantee thorough coverage.
Our findings highlight the importance of refining prompt engineering to fully leverage LLM capabilities across diverse testing scenarios. Improved prompting, combined with exploring SBST methodologies, shows potential for enhancing test suite quality and effectiveness. Additionally, the better performance of GPT-3.5 Turbo over GPT-4 in the Custom dataset requires cautious interpretation. Further analysis is needed to determine if this is due to actual model advancements or the limited sample size.
4.4. [RQ4]: Test Smell detection.
[Experiment Goal]: Our aim is to evaluate the extent to which the test suites generated by LLMs exhibit test smells compared with those produced by an SBST method like EvoSuite.
[Experiment Design]: This experiment assesses the test smell prevalence in test suites generated by LLMs and EvoSuite. CMD was introduced to evaluate the generalizability of LLM-generated test suites on modern and unfamiliar codebases. GPT-4 and Mixtral-8x7B were evaluated only on the CMD due to budget constraints and the focus on understanding the models’ performance on novel data. For EvoSuite, the experiment was applied only to the SF110 and Defects4J, as EvoSuite is incompatible with CMD (cf. Section 3.6). Test smells were detected using TsDetect (Peruma et al., 2020), a state-of-the-art tool that identifies 20 different types of test smells (Peruma et al., 2019, 2020). Due to space constraints, we provide a list of the test smells detectable by TsDetect with their descriptions in our replication package.
The evaluation involved analyzing the test suites for common test smells, such as incomplete assertions, complex test logic, or excessive setup methods. We recorded the frequency and types of test smells detected in each test suite, allowing us to compare the absence of test smells of LLM-generated tests against those produced by EvoSuite.
[Experiment Results]: Table 7 shows that the most common test smells (Van Deursen et al., 2001; Meszaros et al., 2003; Peruma et al., 2019, 2020) produced by both LLMs and EvoSuite are Assertion Roulette (AR), Conditional Logic Test (CLT), Constructor Initialization (CI), Empty Test (EM), Exception Handling (EH), General Fixture (GF), Mystery Guest (MG), and Magic Number Test (MNT), among others. Smells like Default Test, Verbose Test, and Dependent Test were not found in any tests. The results reveal notable differences in test quality across prompt techniques and models, with techniques like COT and GToT producing higher-quality tests, while Few-shot Learning and EvoSuite tend to generate tests with more test smells. The values marked as ”-” for Mistral 7B CMD with ZSL and CoT in Table 7 are due to the absence of compilable test cases for these particular prompt techniques.
For GPT-3.5 Turbo, CoT produces relatively clean tests, with Assertion Roulette in 40.06% of Defects4J files, though Magic Number Test remains high at 99.28%. Few-shot Learning performs worse, with 73.15% affected by Assertion Roulette. On CMD, CoT reduces Assertion Roulette to 6.4%, but Magic Number Test persists at 100%. GToT further reduces Assertion Roulette to 19.86%, though Magic Number Test remains. With GPT-4, the trends are similar, with CoT reducing Assertion Roulette to 14.56% on CMD but Magic Number Test still at 98%. ToT and GToT show slightly lower Assertion Roulette but struggle with Magic Number Test.
Regarding Mistral7B, CoT and Few-shot Learning result in high Assertion Roulette on Defects4J (65.13% and 79.35%), while GToT reduces it to 27.97%. On CMD, GToT performs better, reducing Assertion Roulette to 16.67%, though Magic Number Test persists in 66.67% of files. For Mixtral8x7b, used on CMD, CoT generates few smells, but Magic Number Test remains at 100%. Few-shot Learning fares better with 33.33% showing Assertion Roulette, but Magic Number Test persists. Zero-shot Learning is less effective, with 75% showing Assertion Roulette.
EvoSuite shows similar trends, with 38.24% Assertion Roulette on Defects4J and 100% Magic Number Test. Other significant smells include General Fixture (76.94%) and Redundant Assertion (54.67%) on Defects4J, and higher Assertion Roulette (50.78%) on SF110, along with General Fixture (62.95%) and Lazy Test (6.73%).
| Model | Dataset | Prompt | Test Smell (%) | |||||||||||||||||
| AR | CLT | CI | EM | EH | GF | MG | RP | RA | SE | ST | EA | LT | DA | UT | IT | RO | MNT | |||
| GPT 3 | SF110 | ZSL | 46.17 | 2.68 | 0.16 | 12.71 | 13.02 | 11.13 | 2.92 | 0.0 | 0.79 | 23.52 | 0.39 | 48.38 | 53.75 | 9.31 | 27.70 | 0.0 | 3.24 | 99.84 |
| FSL | 74.44 | 3.91 | 0.11 | 7.14 | 14.51 | 6.70 | 3.57 | 0.0 | 3.13 | 33.15 | 0.45 | 70.65 | 53.91 | 16.52 | 16.29 | 0.0 | 4.13 | 100 | ||
| CoT | 40.98 | 2.15 | 0.0 | 22.25 | 11.60 | 5.93 | 1.98 | 0.0 | 0.95 | 17.53 | 0.09 | 42.44 | 50.26 | 6.53 | 34.54 | 0.0 | 1.89 | 99.05 | ||
| ToT | 59.77 | 3.22 | 0.11 | 4.94 | 19.20 | 13.45 | 2.41 | 0.34 | 3.56 | 28.51 | 0.92 | 56.90 | 76.21 | 13.33 | 15.06 | 0.11 | 2.64 | 100 | ||
| GToT | 38.23 | 0.58 | 0.0 | 28.97 | 10.24 | 2.77 | 1.99 | 0.0 | 2.13 | 11.11 | 0.05 | 42.84 | 40.03 | 3.30 | 37.17 | 0.0 | 2.23 | 100 | ||
| Defects4J | ZSL | 57.01 | 4.05 | 0.06 | 15.72 | 24.30 | 7.50 | 1.09 | 0.0 | 0.91 | 26.90 | 0.79 | 48.55 | 60.70 | 9.92 | 19.65 | 0.0 | 1.09 | 99.64 | |
| FSL | 73.15 | 4.90 | 0.24 | 13.18 | 24.68 | 3.94 | 0.88 | 0.0 | 2.09 | 31.19 | 1.61 | 64.95 | 57.64 | 17.36 | 15.19 | 0.0 | 0.96 | 100 | ||
| CoT | 40.06 | 2.83 | 0.11 | 34.40 | 16.79 | 2.99 | 0.61 | 0.0 | 0.66 | 17.73 | 0.61 | 36.45 | 46.81 | 5.71 | 41.72 | 0.0 | 0.66 | 99.28 | ||
| ToT | 69.56 | 5.52 | 0.32 | 4.57 | 30.60 | 11.20 | 1.03 | 0.0 | 0.79 | 29.81 | 0.79 | 52.92 | 75.95 | 11.20 | 6.23 | 0.0 | 1.0 | 99.97 | ||
| GToT | 29.64 | 2.06 | 0.0 | 52.39 | 12.67 | 2.03 | 0.90 | 0.0 | 1.11 | 7.97 | 0.24 | 28.42 | 32.22 | 1.45 | 55.05 | 0.0 | 1.0 | 99.97 | ||
| CMD | ZSL | 37.72 | 0.0 | 0.0 | 33.33 | 15.79 | 35.96 | 0.0 | 0.0 | 0.0 | 2.63 | 0.0 | 30.70 | 27.19 | 0.0 | 54.39 | 0.0 | 0.0 | 100 | |
| FSL | 18.45 | 0.97 | 0.0 | 49.51 | 25.24 | 14.56 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 18.45 | 16.50 | 0.0 | 77.67 | 0.0 | 0.0 | 100 | ||
| CoT | 6.4 | 0.0 | 0.0 | 78.80 | 7.20 | 2.80 | 0.0 | 0.0 | 0.40 | 0.0 | 0.0 | 6.40 | 7.20 | 0.0 | 89.20 | 0.0 | 0.0 | 100 | ||
| ToT | 15.04 | 0.0 | 0.0 | 71.68 | 7.96 | 18.58 | 0.0 | 0.0 | 0.0 | 0.88 | 0.0 | 17.70 | 15.04 | 0.88 | 76.99 | 0.0 | 0.0 | 100 | ||
| GToT | 19.86 | 0.0 | 0.0 | 58.90 | 15.07 | 17.81 | 0.0 | 0.0 | 0.68 | 1.37 | 0.0 | 18.49 | 21.23 | 0.0 | 74.66 | 0.0 | 0.0 | 100 | ||
| GPT 4 | CMD | ZSL | 30.43 | 1.45 | 0.0 | 11.59 | 14.49 | 20.29 | 0.0 | 0.0 | 1.45 | 2.90 | 0.0 | 36.23 | 36.23 | 0.0 | 34.78 | 0.0 | 0.0 | 98.55 |
| FSL | 41.18 | 1.18 | 0.0 | 10.59 | 31.76 | 20.0 | 0.0 | 1.18 | 0.0 | 2.35 | 0.0 | 48.24 | 45.88 | 2.35 | 36.47 | 0.0 | 0.0 | 100 | ||
| CoT | 14.56 | 0.0 | 0.0 | 28.16 | 18.45 | 17.48 | 0.0 | 0.0 | 0.0 | 0.97 | 0.0 | 18.45 | 35.92 | 0.0 | 54.37 | 0.0 | 0.0 | 98.06 | ||
| ToT | 16.90 | 0.0 | 0.0 | 33.80 | 16.90 | 15.49 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 23.94 | 22.54 | 0.0 | 53.52 | 0.0 | 0.0 | 97.18 | ||
| GToT | 13.98 | 0.0 | 0.0 | 49.46 | 17.20 | 19.35 | 0.0 | 0.0 | 1.08 | 0.0 | 0.0 | 16.13 | 21.51 | 0.0 | 69.89 | 1.08 | 0.0 | 97.85 | ||
| Mistral 7B | SF110 | ZSL | 63.45 | 9.66 | 0.0 | 0.0 | 14.71 | 7.98 | 0.42 | 1.26 | 1.26 | 31.51 | 0.84 | 65.13 | 80.25 | 10.08 | 3.36 | 0.42 | 0.42 | 100 |
| FSL | 72.69 | 11.89 | 0.0 | 0.88 | 7.05 | 0.44 | 0.0 | 1.32 | 6.17 | 27.75 | 0.0 | 68.28 | 69.60 | 15.86 | 5.29 | 0.0 | 0.0 | 100 | ||
| CoT | 52.26 | 7.04 | 0.0 | 2.51 | 19.10 | 2.01 | 0.5 | 1.51 | 1.51 | 22.61 | 0.5 | 53.27 | 84.92 | 11.56 | 5.53 | 0.0 | 0.5 | 94.47 | ||
| ToT | 42.59 | 2.65 | 0.0 | 8.47 | 12.70 | 2.38 | 0.26 | 0.26 | 1.59 | 10.58 | 0.0 | 43.12 | 42.59 | 8.99 | 11.11 | 0.0 | 0.26 | 92.59 | ||
| GToT | 21.20 | 0.82 | 0.14 | 13.54 | 6.57 | 0.82 | 0.14 | 0.0 | 0.55 | 3.42 | 0.0 | 27.91 | 16.01 | 2.74 | 21.34 | 0.0 | 0.14 | 85.64 | ||
| Defects4J | ZSL | 73.64 | 14.34 | 0.39 | 0.0 | 17.44 | 1.55 | 0.78 | 0.78 | 1.55 | 13.18 | 1.55 | 48.06 | 58.14 | 32.17 | 2.33 | 0.0 | 0.39 | 100 | |
| FSL | 79.35 | 14.17 | 0.4 | 1.62 | 8.5 | 0.81 | 0.4 | 0.0 | 1.21 | 25.51 | 0.81 | 62.75 | 78.95 | 21.46 | 2.02 | 0.4 | 0.4 | 100 | ||
| CoT | 65.13 | 13.82 | 0.0 | 0.66 | 21.05 | 0.66 | 0.0 | 0.0 | 1.32 | 10.53 | 1.32 | 41.45 | 81.58 | 31.58 | 0.66 | 0.0 | 0.0 | 95.39 | ||
| ToT | 69.84 | 7.54 | 0.4 | 1.59 | 22.62 | 2.38 | 0.0 | 0.0 | 1.59 | 22.22 | 1.59 | 64.29 | 78.97 | 23.41 | 3.17 | 0.0 | 0.0 | 98.41 | ||
| GToT | 27.97 | 2.89 | 0.0 | 12.86 | 13.67 | 0.0 | 0.0 | 0.0 | 0.16 | 4.82 | 0.16 | 28.94 | 30.55 | 4.18 | 16.08 | 0.0 | 0.0 | 84.41 | ||
| CMD | ZSL | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | |
| FSL | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 100 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 100 | 0.0 | 0.0 | 100 | ||
| CoT | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | ||
| ToT | 0.0 | 0.0 | 0.0 | 0.0 | 100 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 100 | 0.0 | 100 | 0.0 | 0.0 | 100 | ||
| GToT | 16.67 | 0.0 | 0.0 | 33.33 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 16.67 | 0.0 | 0.0 | 33.33 | 0.0 | 0.0 | 66.67 | ||
| Mixtral8x7B | CMD | ZSL | 75 | 0.0 | 0.0 | 0.0 | 25.0 | 75.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 25.0 | 25.0 | 0.0 | 0.0 | 0.0 | 0.0 | 100 |
| FSL | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 33.33 | 0.0 | 0.0 | 33.33 | 0.0 | 0.0 | 66.67 | ||
| CoT | 0.0 | 0.0 | 0.0 | 0.0 | 50 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 50.0 | 0.0 | 50.0 | 0.0 | 0.0 | 100 | ||
| ToT | 44.44 | 11.11 | 0.0 | 44.44 | 11.11 | 0.0 | 0.0 | 0.0 | 0.0 | 11.11 | 0.0 | 55.56 | 33.33 | 0.0 | 55.56 | 0.0 | 0.0 | 100 | ||
| GToT | 0.0 | 0.0 | 0.0 | 50.0 | 50.0 | 50.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 50.0 | 0.0 | 100 | 0.0 | 0.0 | 100 | ||
| EvoSuite | SF110 | 50.78 | 0.0 | 0.0 | 6.90 | 93.10 | 0.0 | 4.44 | 0.0 | 0.0 | 17.53 | 0.0 | 54.67 | 76.94 | 0.07 | 6.90 | 0.0 | 3.96 | 100 | |
| Defects4J | 38.24 | 0.0 | 0.0 | 6.90 | 93.10 | 0.0 | 4.44 | 0.0 | 0.0 | 10.71 | 0.0 | 62.95 | 82.18 | 6.73 | 0.38 | 0.0 | 4.34 | 100 | ||
1
∗ AR: Assertion Roulette, CLT: Conditional Logic Test, CI: Constructor Initialization, EM: Empty Test, EH: Exception Handling, GF: General Fixture, MG: Mistery Guest, RP: Redundant Print, RA: Redundant Assertion, SE: Sensitive Equality, ST: Sleepy Test, EA: Eager Test, LT: Lazy Test, DA: Duplicate Assert, UT: Unknown Test, IT: Ignored Test, RO: Resource Optimism, MNT: Magic Number Test, ZSL: Zero-shot Learning, FSL: Few-shot Learning, CoT: Chain-of-Thought, ToT: Tree-of-Thoughts, GToT: Guided Tree-of-Thoughts.
5. Threats to Validity
This section addresses the threats to the validity of this study.
External Validity. The generalizability of our findings to real-world software testing may be limited by several factors. Our benchmarks (Defects4J, SF110, and CMD) may not fully capture the complexities of real-world software. The CMD dataset, with only two projects, limits diversity and representativeness. This small sample size may affect results, particularly for code coverage, where GPT-3.5 Turbo performed better than GPT-4 on CMD—a result that requires cautious interpretation. Further analysis across larger datasets is needed to confirm trends. Additionally, overestimation of LLM capabilities could occur if models were trained on test cases within the datasets, though we mitigated this with CMD. Focusing on four instruct-tuned LLMs from two families (OpenAI and MistralAI) also limits representativeness. Furthermore, the knowledge cutoff dates for Mistral models may affect performance, as Mistral 7B has a 2021-2022 cutoff, and Mixtral 8x7B was released in December 2023. Our findings reflect the models’ capabilities during the study period (November 2023 to April 2024).
Internal Validity. Internal threats stem from the randomness inherent in LLM outputs and EvoSuite, prompt design, and the test smell detection methodology. To account for randomness, we repeated each experiment 30 times and applied statistical analysis to ensure reliable interpretations. While our prompts were designed based on model guidelines and prior research, they may not be fully optimized for every model or task. Our test smell detection methodology, although comprehensive, may miss other quality issues such as performance or security concerns within the test suites. We also need to consider flaky tests, which can result in inconsistent test smell detection outcomes.
6. Related work
Siddiq et al. (Siddiq et al., 2023) evaluated LLMs like Codex, GPT-3.5-Turbo, and StarCoder for generating JUnit tests, focusing on compilation rates, correctness, coverage, and test smells. Our study aligns with their findings on challenges like persistent test smells, including Magic Number Test and Assertion Roulette. However, we further explored the impact of various prompt techniques to improve test quality and reduce smells. Additionally, we assessed a broader range of LLMs and prompt strategies, providing a more comprehensive analysis of test quality.
Tang et al. (Tang et al., 2024) evaluated ChatGPT and EvoSuite for generating Java unit tests, focusing on correctness, readability, and coverage. Like our study, they found that ChatGPT underperforms in coverage compared to EvoSuite, though it showed potential in generating diverse test inputs. EvoSuite consistently achieved higher coverage overall. Our study extends this by assessing multiple prompt techniques and focusing on test smell detection, offering a deeper understanding of the maintainability and quality of LLM-generated tests.
Alshahwan et al. (Alshahwan et al., 2024) introduced TestGen-LLM, a tool at Meta that improves existing human-written tests using LLMs. Unlike our study, which focuses on generating tests from scratch with various prompts, TestGen-LLM enhances tests by increasing coverage without introducing regressions. During Meta’s test-a-thons, the tool had a 73% acceptance rate for production deployment. While we focus on prompt engineering and test smell detection, both approaches aim to generate high-quality tests using LLMs.
7. Discussion of Findings and Implications
Our findings indicate several key implications for both researchers and developers. The prompt technique plays a crucial role in the quality of LLM-generated test suites. CoT and GToT generally produce better results, indicating that refined prompt designs lead to improved outputs. While LLMs like GPT-3.5 and GPT-4 show potential, especially with CoT, they still lag behind EvoSuite in terms of coverage. However, LLMs excel in generating more readable tests when using well-crafted prompts, suggesting they could complement tools like EvoSuite in scenarios prioritizing readability. A key issue is the persistence of test smells like Magic Number Test across all models and techniques, highlighting the need for human intervention to improve test maintainability. GPT-3.5 Turbo’s better performance over GPT-4 on CMD suggests differences in handling novel codebases, though CMD’s small size calls for cautious interpretation, with larger datasets needed for validation.
8. Conclusion
We conducted a large-scale study on the effectiveness of LLMs for unit test generation. Our experimental results show, for the first time in the literature, the substantial impact of prompt design on the quality of LLM-generated unit tests. Techniques such as Chain-of-Thought (CoT) and our newly introduced Guided Tree-of-Thought (GToT) prompting strategy produced the highest-quality tests. Although correctness issues persist, compilable LLM-generated unit tests closely resembled human-written ones in terms of adherence to coding standards, highlighting their potential. A key trade-off emerged: while EvoSuite consistently achieved higher code coverage, CoT-prompted GPT-3.5 Turbo and Mistral 7B excelled in generating more readable and maintainable tests. However, these advantages were primarily observed on familiar datasets, pointing to potential limitations in handling unfamiliar codebases. These preliminary, yet promising, results suggest that further research is needed to fully unlock the potential of LLMs in unit test generation, particularly in improving test quality by addressing persistent issues like test smells.
9. Data Availability
To promote transparency and facilitate reproducibility, we make our artifacts available to the community at:
https://anonymous.4open.science/r/LLM4TS-0F76/
The repository includes all LLM and SBST-generated tests, our Custome dataset, and all scripts for the replication of our study.